A Complete Bibliography of Publications in ACM SIGAPP Applied Computing Review

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

20 August 2024 Version 1.05

Title word cross-reference

k [BBH11].

11g [MMC12]. 12000 [LLMW93].

2001 [Dam01b, FS02, vD02]. 23rd [FS02].

3d [CYLL16]. 3d-stacked [CYLL16].

6 [UBA^{+}23].

802.15.4 [TFSO14]. 802.15.4-based [TFSO14].

90 [NDS96]. abstracting [KR13]. Accelerated [HEK^{+}13]. Accelerating [GG11]. Acceleration [LWS^{+}23]. access [MMSV02, SSR21]. accessibility [MCP15]. accessible [OTYM01]. accurate [HAG21, WLK^{+}21]. achievements [Dij99]. activation [GD14]. Active [BHK^{+}15, BDMT16, QHW^{+}13, Raj99a]. Actor [AAR^{+}24]. adaptation [DOP^{+}14, GLPT12]. Adapting [BBSB24]. Adaptive [PPM18, AC20, ARS15, FF11, PEVR13, TH19, WZL^{+}22]. adaptors [ATS96]. Addressing [YPB19]. adjusting [CYLL16]. Advancing [ÇÖT24, DQG^{+}23]. advantage [War95]. Adversarial [LOCD23]. advertisement [AJS^{+}13]. advisory [SKH93]. affect [AJS^{+}13]. after [LLC16]. Agent [SPB01, YLJ24, Ano01b, VF01]. agents [BWLP01, Det01, JL99, WW01]. aggregation [RHA11]. Agile [SAG23, FS02]. AI [PVT24]. aided [PSS17]. aircraft [TO94]. alarm [VFMM94]. alert [HF19]. algorithm [KHC^{+}11, KL17, LHO12, Rai99, Ros19, SPPP11]. algorithms [HK15, SP16, WC20]. Alignment [Hua93]. allocation [DCUA20]. amplicon [BP18]. Analysing [SB96]. Analysis [LHO12, PeSNY23, ABLRC21, FMNP94, LCM18, LS21, PKBH17, SD12, SD13, SS02, TJ12, VSF97, VHS20, VGH^{+}16, ZCG^{+}16]. analytical [DOP^{+}14]. Analyzing [BDWP20]. Android [MCMS22, ZCG^{+}16]. aneurysm [LBKV01]. annotations [MCY^{+}19]. anomalies [BP18, PC19]. anomaly [FF11, PSS17]. anonymity [AHH12]. Answering [FT19]. applets [CR98]. application [HURU11, LLMW93, LSV12, LCL12, RFB^{+}15, SS93, SW99]. applications [ARS15, CPT10, GWB20, HYL^{+}17, Jan93, KP18, LCM18, LS21, LF19, MBR17, PZ18]. applied [TAdOD13]. Applying [CRX18, LSPC23]. Approach [MA22, AJS^{+}13, ABLRC21, BS97, HMY95, HK17, HDC^{+}14, Ish96, KKEK17, KR13, MZAS14, MBR17, NL21, Rai99, SSR21, SS93, SHV13, SD12, SER15, SB16, SKH93, SL99, YSCX17]. Approaches [AAR^{+}24]. architectural [DAI^{+}21]. Architecture [SH23, CDT18, Cit93, LW12, PZ18, VF01, YKdSM14, vD02]. architectures [GRJD15, KKEK17]. archival [HMY95]. area [SWS20]. arrays [JPM16]. artificial [SD12]. Aspect [CCDS13, LFT23, HK17]. Aspect-Based [LFT23, HK17]. Aspect-driven [CCDS13]. AspectJ [RZW01]. Assessment [SGPM22]. assignment [KL15]. assistance [VFMM94]. assistants [NYKI98]. assisted [Tre02, XDS22]. Association [LWS^{+}23, AB12, MN18]. Astrophysics [PeSNY23]. attack [SSE18]. attacks [HF19]. attention [LNLC16, XDS22]. attribute [BHK^{+}15]. attributes [PIR20]. authoring [MVOdGCP11]. AutoFix [YSCX17]. Automated [CC17, FRNNS18, YSCX17]. Automatic [MVOdGCP11, PVT24, LC13]. Automating [OSG23]. automotive [SB16]. autonomous [KCDS98, PKVM21, STP^{+}12, SSS^{+}11]. avoidance [SSS^{+}11]. avoiding [TFSO14]. Aware [BZT23, BBSB24, CCDS13, HCC^{+}17, KJMJ10, TK22, WLK^{+}21]. back [Alo18]. bandwidth [ZPH^{+}15]. Based [AKBJ14, LFT23, LWS^{+}23, MA22, SH23, AHH12, Alo18, AdSR^{+}12, BHK^{+}15, BDMT16, CDAR11, CO00, DF96, DCUA20, FRNNS18, FF11, GRJD15, HF19, HW24, HK17, HEK^{+}13, HYL^{+}17, HCC^{+}15, HCT^{+}13, KP18, KTH18, LWC^{+}12, MBR17, MCP15, MYX^{+}14, NL21, NHM15, PND^{+}13, QM15, QHW^{+}13, RFB^{+}15, SSR21, SHV13, SD13, SMK17, SSS^{+}11, SSQJ99, SL99, TJ12, THW14, Tre02, TFSO14, UOB^{+}13, VPB19, VSF97, Van02, WdCS21, YHTN01, YKdSM14, YAG^{+}15, ZCG^{+}16]. behavior [BK19, MCY^{+}19, MR01, SPB01]. behavior-driven [BK19]. Behavioral [LOCD23, RP11]. Belief [Baa23]. Benchmarking [GWB20]. benchmarks [WCS20]. Best [PVT24, Dam01b]. between [TD14]. bigraphs [GRJD15]. Binary [PVT24, HLF^{+}19]. Binning [PeSNY23]. Biologically [AB12]. black [NVP21]. black-box [NVP21]. blackboard [SS93]. blended [MAM^{+}16]. Block [BZT23]. Block-RACS [BZT23]. blockchain [SGPM22]. Blocking [MA22, LC13]. blogosphere [HK15]. body [SWS20]. boosted [PSS17, SD13]. Botnets [LOCD23]. bounded [LCL12]. box [NVP21]. BPEL [CPT10]. Brazilian [MAM^{+}16]. browsing [MCY^{+}19]. Buffer [SHV13, FD15]. bug [IZ21]. bug-fixing [IZ21]. buildings [LMLAK17]. built [LSV12]. bulk [MZ97]. business [CC17, KR13, PC19]. BYOD [SSE18]. bypass [CCSK11, VGM^{+}20]. Byzantine [KRHH21].

C [BP18, Chi96, SHV13, SW99, YSCX17]. C/C [SW99]. cache [HCT^{+}13, LCC^{+}21]. caching [SPPP11, WZL^{+}22, YHTN01]. Cagliari [FS02]. calculus [CW19]. call [GDJ96]. cancer [MN18]. cardinality [Baa19]. cars [PKVM21]. Case [RZW01, BC94, MAM^{+}16, VGM^{+}20]. Catalog [ÇÖT24]. Catalogue [PVT24]. Categorizing [SAG23]. CC [Kes96]. CD [HMY95]. CD-ROM [HMY95]. CD-ROM/WWW [HMY95]. CEMDA [MCMS22]. centers [GSKJ18, VFMM94]. centres [NOVS11]. cGA [TAdOD13]. Challenges [SAG23, Dij99, RC99]. Change [Baa23, IZ21]. changes [IZ21]. channel [LF19]. character [ABCW95]. characterization [AW95]. characterizing [CGR99]. checking [FRNNS18]. Chief [Ber94]. CHIEv [LS21]. choice [Raj99b]. Cholesky [Alo18]. Circular [SHP22]. Classification [LFT23, AB12, BS97, DF96, HCC^{+}15, MBP12, SUV97, WdCS21]. classification-based [HCC^{+}15]. classifier [SD13]. classifiers [TD14]. Client [BZT23]. clinical [AGH^{+}21]. clone [MRS12, WC20, WCS20, ZSRS13]. clothing [Ros98, Ros19]. cloud [LK13, WC20, ZCG^{+}16]. cloud-based [ZCG^{+}16]. clustering [BAK22, DCUA20, HYL^{+}17, HMB12]. clustering-based [DCUA20, HYL^{+}17]. CMPs [CYLL16]. Co [KK97, FS02, LF19]. co-located [FS02]. Co-operative [KK97]. co-resident [LF19]. Code [SH23, DF96, DAI^{+}21, DF97, LLS12, WC20, WCS20]. code-clone [WC20, WCS20]. Code-to-Architecture [SH23]. coded [JPM16]. coding [MN18]. codings [Rai99]. coexistence [OWB^{+}14]. collaboration [OWB^{+}14]. Collaborative [GAB^{+}23, OTYM01, BHK^{+}15, FS12, LWC^{+}12, SSW02]. collections [Raj99a]. Combining [BCI18]. Communication [AKBJ14, MYX^{+}14]. communications [Cit93, HYL^{+}17, HXGB22]. Community [NYKI98, UOB^{+}13]. Companies [SAG23]. company [Rin00]. Comparative [WA94]. competition [SPD^{+}19]. Competitive [War95]. compilation [DL16]. completion [XDS22, YLY^{+}22]. Complex [MCMS22, MBR13]. components [BS97, CP14, DF96, GD14, VWT13]. Composite [FdE11, TJ12]. compositional [Kes96]. compositions [BBH11]. Comprehensive [ÇÖT24]. compressed [MPFC12]. Compression [HW24]. computation [CDAR11, Kes96]. computations [Bad96, GG11]. Computer [PSS17, LLMW93, LDY20, Tre02]. computer-assisted [Tre02]. Computing [Dij99, LWS^{+}23, SN96, Chi96, HCC^{+}15, Ish96, JCP17, Muk99, Nic02, RC99, Raj99b, Raj99c, Vil99, WC20]. concavities [DM21]. concepts [Baa19]. concurrency [OKK13]. Concurrent [Jan93, CW19, HAG21, LS21, RZW01]. conference [FS02]. configurable [LCM18]. conflict [FS12]. Consensus [YLJ24]. considerations [KL15]. consistency [FRNNS18]. consolidation [LCC^{+}21]. constraint [AM98]. consumer [FD15]. consumers [BSG16]. consumption [VWT13]. container [Rai99]. content [AdSR^{+}12]. content-based [AdSR^{+}12]. context [CCDS13, PEVR13, WZL^{+}22]. context-aware [CCDS13]. Contextual [CDT18]. Contract [BEL24, WCR95]. Contraction [Baa23]. Contrastively [MA22]. Control [AKBJ14, AAR^{+}24, YLJ24, Ano01b, BWLP01, CYLL16, Muk99, OKK13, SSR21, TBC94, VFMM94]. controller [LW12]. controllers [TO94]. Convolutional [HW24]. Cooperative [AKBJ14, AM98]. Coordinating [CR98, KCDS98, NOVS11]. Coordination [AM98, UOB^{+}13]. core [SWH^{+}13]. cores [AC20, CYLL16]. coronavirus [APPV21]. correlated [YHTN01]. Correlation [DCUA20, HF19, NHM15, Nyg94]. Cost [SSQJ99, MPSR00, WCR95, WZL^{+}22]. Cost-based [SSQJ99]. cost-efficient [WZL^{+}22]. courses [MAM^{+}16]. CPU [LWS^{+}23, VWT13]. CPU/FPGA [LWS^{+}23]. crawling [LS21]. create [CMDM20]. creation [Tre02]. creational [CP14]. Critic [AAR^{+}24]. critical [FRNNS18, NRH13]. Cross [LFT23, HMY95, SER15, TCPC17, UOB^{+}13]. cross-community [UOB^{+}13]. Cross-Domain [LFT23]. cross-layer [TCPC17]. cross-platform [HMY95]. crowdsensing [PZ18]. crowdsourcing [LMLAK17]. CSPs [LY13]. current [CDT18]. customer [NL21]. cyber [BDWP20, LCC^{+}21]. cyber-physical [BDWP20, LCC^{+}21]. Cytochrome [BP18].

D2D [HXGB22]. DAG [QM15]. Data [DQG^{+}23, HW24, OKK13, PeSNY23, VHS20, Alo18, AB12, CGP02, CCDS13, FT19, FRS16, GSKJ18, KRHH21, LCM18, LCL12, MZ97, PKBH17, PND^{+}13, RHA11, TJ12, WHC19, YHTN01, YPB19]. data-flow [TJ12]. data-reflective [CCDS13]. databases [GG11, Han95, WHC19]. Datasets [DFP^{+}22, BAK22]. Deadline [BBSB24]. Deadline-Aware [BBSB24]. Dealing [CO00, MMC12]. Debt [ÇÖT24, DAI^{+}21]. deceptive [LHO12]. decision [AGH^{+}21, WCR95]. declarative [LSV12]. DECmpp [LLMW93]. decomposing [Ros98, Ros19]. decomposition [Alo18]. deduplication [LF19]. Deep [DQG^{+}23, HW24, WdCS21, BCI18, WLK^{+}21, WJdCS22]. defense [THW14]. defined [CRX18]. degradation [DAI^{+}21]. delay [CRX18]. demise [Ber95a]. demonstration [SSW02]. Dependencies [GAB^{+}23]. deployed [HAG21]. Deployment [GD14, GRJD15]. derivation [CGP02]. description [Baa19]. Design [BBSB24, DF96, DQG^{+}23, PZ18, WJdCS22, CCDS13, CP14, FMNP94, Gea94, HCT^{+}13, LCC^{+}21, LDY20, NL21, NRH13, TO94, VF01]. Designing [RC99, RP11]. Detailed [NVP21]. details [Dam01b, Raj99c]. detect [APSHB15]. Detecting [BP18, PC19, TD14]. Detection [BEL24, AJS^{+}13, HF19, KHC^{+}11, LC13, LLS12, LNLC16, LF19, MPNdL10, PSS17, SHP22, SPD^{+}19, WC20, WCS20]. detector [FF11]. develop [SW99]. Developing [PVT24]. Development [SAG23, AM00, CPT10, KKH^{+}17, LLMW93, NRH13, VSF97]. deviations [CO00]. device [HYL^{+}17]. device-to-device [HYL^{+}17]. Devices [BBSB24, AJS^{+}13, DL16, HCT^{+}13]. DGAs [SPD^{+}19]. diagnosis [PSS17]. dialect [Chi96]. dictionaries [MPFC12]. Differential [AAR^{+}24, LOCD23]. digital [LDY20, NYKI98]. Dimension [HW24]. dimensionality [Alo18]. directions [CDT18]. Discovery [DQG^{+}23]. Discriminating [FRS16]. discussion [vD02]. distance [YKdSM14]. Distributed [Muk99, AM98, BHK^{+}15, CR98, CGR99, JL99, Kes96, NVP21, OTYM01, RC99, Raj99b, Raj99c, RZW01, WW01]. distribution [PPM18]. DIWS [LFT23]. DIWS-LCR-Rot-hop [LFT23]. DNA [BAK22]. Document [VPB19, HMB12, JCP17, SD12]. Document-based [VPB19]. document-level [SD12]. documentation [CC17]. documents [APSHB15, MMSV02, MVOdGCP11, Tre02]. DOF [UBA^{+}23]. Domain [LFT23, PeSNY23, VSF97, BSV97, DF97, LY13, MZ97, PZBdBC18, SB16]. Domain-Independent [LFT23]. domain-specific [PZBdBC18, SB16]. DRAMs [CYLL16]. drift [YPB19]. drive [JPM16]. driven [BK19, CCDS13, CP14, LCL12, MBR17]. drives [LW12]. Drop [HXGB22]. Drug [DQG^{+}23]. dual [HMB12]. duplicate [LC13]. DVS [KL15]. Dynamic [SP16, SS02, EHBS21, HLF^{+}19, LCL12, PPM18, WZL^{+}22, YH16]. dynamical [TCPC17]. Dynamics [AAR^{+}24, LNLC16].

e-health [NOVS11]. e-markets [MPNdL10]. E/E [SB16]. Early [DGS^{+}14]. Ecxpert [Nyg94]. EDF [AC20]. Edge [BBSB24]. editor [Raj99b, Ber94]. Editor-in-Chief [Ber94]. editorial [Dam01a, DEM02, Par01]. education [OTYM01]. educational [Van02]. effect [SHP22]. effective [MN18, TFSO14]. effectiveness [MYX^{+}14]. Effects [LOCD23, MCY^{+}19, Val97]. efficiency [BSVV96, KL15, LCC^{+}21, SAL19]. Efficient [BBSB24, CCSK11, DL16, HCC^{+}15, SPPP11, CYLL16, KTH18, KRHH21, KHC^{+}11, LKH17, SM22, SWS20, TH19, WZL^{+}22]. Elastic [PeSNY23]. elasticity [KWBE20]. elderly [MZAS14]. elementary [SER15]. Eliminating [LY13]. emails [TD14]. Embedded [BBSB24, DL16, HCC^{+}15, LKP12]. Embedded/Edge [BBSB24]. embedding [NHM15]. Embeddings [MA22, PIR20]. EMF [WdCS21]. EMF-related [WdCS21]. empirical [IZ21, MRS12]. end [CRX18, HDC^{+}14]. end-to-end [CRX18, HDC^{+}14]. Energy [AKBJ14, BBSB24, KL15, LCC^{+}21, SAL19, SM22, YAG^{+}15, ZPH^{+}15]. energy-and-bandwidth [ZPH^{+}15]. Engineering [Baa23, Ano01a, FS02, SSW02, SS02, VGH^{+}16]. enhancement [LCC^{+}21, MR01]. Enhancing [HLF^{+}19]. ensemble [SD13]. enterprise [CC17, SSE18]. entities [KCDS98]. Entity [MA22, APSHB15, BDMT16, RT20, YA22]. environment [Ang95, Ish96, KGKK96, LLMW93, LCL12, OTYM01, RC99]. environmental [OTYM01]. environments [FS12, SSE18, SSS^{+}11, YH16]. epidemiological [APPV21]. epistatic [LHO12]. Equations [AAR^{+}24]. erasure [JPM16]. erasure-coded [JPM16]. Error [AKBJ14, OSG23, FD12]. Establishing [BWLP01]. Estimating [WZL^{+}22]. Estimation [AAR^{+}24, UBA^{+}23, MPSR00, VWT13, YH16, YKdSM14]. Ethernet [OD93]. Europa [GDJ96]. Evaluate [DFP^{+}22]. Evaluating [CMDM20]. evaluation [CBV97, HK15, KMR20, PZ18, PZBdBC18, TWK^{+}13]. evaluations [QM15]. evasion [SPD^{+}19]. Event [YLJ24, Nyg94]. Event-Triggered [YLJ24]. Events [MCMS22]. everlasting [SPD^{+}19]. Evolutionary [MR01, CDAR11, SP16]. exchange [UOB^{+}13]. execution [HLF^{+}19]. existing [WC20]. experiences [MMSV02]. Experiencing [GS96]. Experiment [LDY20]. experimental [TWK^{+}13]. Expert [LK93, TBC94, VFMM94]. explain [FD12]. explainable [AGH^{+}21]. exploitation [AF21]. Exploiting [RT20, Nyg94]. exploratory [ABLRC21]. expression [FRS16, LNLC16]. expressions [APSHB15]. Expressive [Baa19]. exsisting [DF97]. extended [LV18]. extensibility [CW19]. extensible [LSV12]. extension [MMC12]. extensions [OWB^{+}14]. extent [MMC12]. external [KP18]. Extracting [DF97, HK17]. extraction [BDMT16, CC17, FD12, SL99, WJdCS22]. extractive [GN22]. eXtreme [FS02]. eye [LNLC16].

FaaS [KWBE20]. face [NHM15]. faceted [BS97]. facial [LNLC16]. factoring [BVC^{+}14]. Factors [LOCD23, KKH^{+}17]. facts [ZSRS13]. Fairness [SWH^{+}13]. fall [MZAS14]. fault [KHC^{+}11, RM12]. faulty [GD14]. feasibility [LCM18]. Feature [AdSR^{+}12, ABLRC21, TD14, YPB19]. Federated [BZT23]. feedback [LMLAK17, VPB19]. few [YLY^{+}22]. few-shot [YLY^{+}22]. fictions [ZSRS13]. Fiducial [UBA^{+}23, LKH17]. field [KTH18, LBKV01, TH19]. file [HCT^{+}13, KJMJ10]. Files [MCMS22]. filtering [LWC^{+}12]. Fine [YLY^{+}22, WJdCS22]. Fine-grained [YLY^{+}22]. fine-tuned [WJdCS22]. Finite [YLJ24]. Finite-Time [YLJ24]. fixing [IZ21, YSCX17]. flash [HCT^{+}13, KHC^{+}11, PND^{+}13]. flash-based [PND^{+}13]. flexible [KR13, KGKK96, LK13, LV18, NRH13]. flow [LCM18, LBKV01, TJ12, TCPC17, YSCX17]. flow-oriented [LCM18]. fly [TO94]. fly-by-wire [TO94]. focus [PZBdBC18]. forecasting [YAG^{+}15]. foreign [MMSV02]. formal [DF97, GRJD15, MBR17, SUV97]. formation [NYKI98]. Fortran [NDS96]. forum [vD02]. FPGA [LWS^{+}23]. Framework [DFP^{+}22, AHH12, BBH11, CBV97, EHBS21, HMB12, KP18, KTH18, KKH^{+}17, LSV12, MBR13, MZ97, NRH13, PZBdBC18, Raj99a, Tre02, TH19, UOB^{+}13, VSF97]. framework-based [VSF97]. frameworks [Val97]. Fraud [MPNdL10, VGM^{+}20]. free [FD15]. Frequency [AAR^{+}24, CYLL16]. Full [UBA^{+}23]. Functional [LCM18, SER15]. future [CDT18, Nic02, Raj99b, WCS20]. fuzzy [BS97, BBH11, CO00, DF96, MA00, MKR93, MPSR00, PEVR13].

game [THW14]. games [DGS^{+}14]. Gamification [LSPC23]. gateway [SW16]. Gaussian [BBSB24]. gaze [LNLC16]. Gender [PVT24]. gene [FRS16]. Genealogical [ZSRS13]. Generating [GN22, SH23]. generation [LWK14, TH19]. generative [DF97]. generic [HDC^{+}14]. genetic [BDMT16, LHO12, Rai99, VHS20]. Genome [LWS^{+}23]. Genome-Wide [LWS^{+}23]. genre [HK17]. geographic [CCSK11]. global [AC20, KKH^{+}17, RC99]. global-scale [RC99]. GNU [ZPH^{+}15]. goal [BDWP20]. government [LK93]. grained [YLY^{+}22]. grammars [BHK^{+}15]. Graph [MA22, FRS16, HF19, RT20, XDS22, YLY^{+}22]. Graph-Based [MA22, HF19]. graphical [BK19]. graphs [DCUA20]. greedy [Ros19]. green [LWK14]. grid [MYX^{+}14, YAG^{+}15]. group [PZBdBC18]. grouping [TFSO14, TCPC17]. groups [TK22]. guaranteed [FD12]. Guaranteeing [HDC^{+}14]. guarded [BHK^{+}15]. Guest [Dam01a, DEM02, Par01, Raj99b]. guide [War95].

Hammer [HXGB22]. hard [QM15]. Hardware [OWB^{+}14]. harmonizing [RFB^{+}15]. Harnessing [DQG^{+}23]. hashing [DOP^{+}14]. Hazard [BC94]. health [NOVS11, UOB^{+}13]. HEAT [SM22]. heavy [VGM^{+}20]. heavy-hitters [VGM^{+}20]. help [APSHB15]. Hensel [LLS12]. Heterogeneous [LWS^{+}23, AM98, JPM16, KTH18, KL15, SM22, SWH^{+}13]. heuristic [SER15]. hibernation [HCC^{+}15]. hidden [TFSO14]. High [Kes96, Chi96, LW12, WCR95]. high-performance [LW12]. high-yield [WCR95]. highly [LCM18]. highly-configurable [LCM18]. Hitec [GS96]. hitters [VGM^{+}20]. hole [CCSK11]. homogeneous [JPM16]. homology [SHP22]. hop [LFT23]. hot [PND^{+}13]. Hough [FF11]. Hough-transform-based [FF11]. HPC [GDJ96]. HTML [Bou95]. human [KKH^{+}17, LNLC16]. humanities [Nic02]. Hybrid [Han95, HCC^{+}17, HCT^{+}13, LS21, PC19, SPPP11, TAdOD13, WHC19]. hyperspace [DOP^{+}14].

I/O [BBSB24]. ICC [Chi96]. identification [BC94, MN18, PND^{+}13]. Identifying [LOCD23, SAG23]. IEEE [TFSO14]. IHE [UOB^{+}13]. image [AdSR^{+}12, MR01, NHM15, VHS20]. images [Lea98]. impact [HURU11]. implementation [WJdCS22]. implementations [HEK^{+}13]. improve [MAM^{+}16, PEVR13]. improvement [Ben00, KKH^{+}17]. Improving [Ano01a, ABLRC21, JPM16, MA00, LLC16]. Impulsive [YLJ24]. incast [TCPC17]. Inclusive [PVT24]. incorporate [NL21]. Increasing [Ano01b, RM12]. Independent [LFT23]. Indicator [ÇÖT24]. Industrial [Ben00, FRNNS18]. industry [KMR20]. inevitable [Ber95a]. Inference [GAB^{+}23]. inferential [Han95]. Inferring [PIR20, LWC^{+}12]. information [ABLRC21, CC17, Det01, Jan93, PPM18, SMK17, WJdCS22, YHTN01]. infrastructure [Bad96, WC20]. InMap [SH23]. Insights [SAG23, ZSRS13]. Integrated [DFP^{+}22, FMNP94]. intelligent [VF01, WCR95]. intensive [LCM18]. inter [SMK17]. inter-stroke [SMK17]. Interaction [TD14]. interactions [FD12]. interactive [MVOdGCP11, MCY^{+}19]. interconnect [VGM^{+}20]. interest [BSG16]. interface [BK19]. interfaces [CCDS13, RP11]. internal [KP18]. internal-external [KP18]. international [FS02]. Internet [AW95, LWK14]. Interpolating [DM21]. interpolation [CMDM20]. interpretation [PEVR13]. interval [FF11]. intra [SMK17]. intra-stroke [SMK17]. Intrinsic [HW24]. introduction [GS96, LLMW93, Raj99b]. inversion [JCP17]. Investigating [MCP15]. investment [SB96]. IoT [GAB^{+}23, SW16]. IP [LU95]. IP4 [TWK^{+}13]. IPV6 [TWK^{+}13]. IR [HAG21]. IR-level [HAG21]. iris [HURU11]. ISO [RFB^{+}15]. ISO-based [RFB^{+}15]. Issue [SN96, Dam01b, Raj99c, Val97]. issues [LU95, RC99, SGPM22]. Italian [MMSV02]. Italy [FS02]. iterative [YPB19]. iterators [TJ12]. IVI [TWK^{+}13].

Java [Tre02, Vil99]. Java-based [Tre02]. JavaCC [SW99]. job [LV18]. join [GG11]. Joint [ZPH^{+}15, AF21].

K12 [Men95]. kernel [OWB^{+}14]. key [LKP12]. Klaim [GLPT12]. Knowledge [CGP02, BCI18, FT19, KK97, PHJJ15, RT20, YLY^{+}22].

LAN [OD93]. landscapes [LHO12]. Language [BV99, BEL24, GN22, PZBdBC18, RFB^{+}15]. Large [BEL24, JCP17, LBKV01, SAG23, Alo18, GG11, KR13, Ros19]. Large-Scale [SAG23, LBKV01]. larger [TCPC17]. larger-scale [TCPC17]. Latent [APPV21]. layer [TCPC17]. lays [Ros98, Ros19]. LCR [LFT23]. LDA [AF21]. LDA-LFM [AF21]. Leading [LOCD23]. leak [YSCX17]. Learned [MA22]. Learning [AAR^{+}24, BZT23, DQG^{+}23, SH23, ABLRC21, BDMT16, BCI18, HXGB22, MN18, PSS17, TD14, WdCS21, WJdCS22, YH16, YAG^{+}15, YLY^{+}22]. Learning-Based [SH23, WdCS21]. legislative [MMSV02]. level [HAG21, RP11, SD12]. levels [CYLL16]. Leveraging [GAB^{+}23]. lexicons [SD12]. LFM [AF21]. libraries [BSV^{+}00]. Library [OSG23]. life [AGH^{+}21, Suc99]. lifetime [WZL^{+}22]. light [KTH18, TH19]. line [PND^{+}13]. linear [NHM15, VPB19]. link [FT19, MH11]. link-translating [MH11]. Linux [SAL19]. literature [WdCS21, WJdCS22]. LLM [BEL24]. local [NHM15]. Localisation [UBA^{+}23, LKH17]. locally [NHM15]. located [FS02]. location [AHH12, TK22]. log [FS02]. logic [Baa19, CO00, LDY20, MA00, SKH93]. logical [WA94]. logistic [MPNdL10]. logs [PC19]. long [MN18]. look [LK93]. Low [BBSB24, DQG^{+}23, WCR95]. low-cost [WCR95]. Low-Data [DQG^{+}23].

Machine [AAR^{+}24, SH23, ABLRC21, GSKJ18, HEK^{+}13, HXGB22, MN18, TD14, YAG^{+}15]. machines [LF19, MBP12, SWH^{+}13, VWT13]. main [SPPP11]. malware [SPD^{+}19, ZCG^{+}16]. Management [ÇÖT24, CGP02, Gea94, HCC^{+}17, HXGB22, KP18, LCL12, LKP12, LK13, SM22, Van02, WHC19]. manufacturing [KGKK96]. Mappers [SH23]. mapping [DAI^{+}21]. Mappings [SH23, FdE11]. maps [MCP15]. Marker [UBA^{+}23]. markers [DGS^{+}14]. market [NL21]. markets [MPNdL10, SPB01]. Maslow [HXGB22]. mass [NL21]. mass-market [NL21]. massive [BAK22]. massively [LLMW93]. Matching [MA22, FdE11]. May [FS02]. means [MMSV02]. Mechanism [BZT23, ATS96, PND^{+}13, SSE18, SUV97]. mechanisms [GD14]. media [HK17, MVOdGCP11]. media-oriented [MVOdGCP11]. memory [HCC^{+}17, HLF^{+}19, KHC^{+}11, LF19, SPPP11, YSCX17]. Mental [MCMS22]. merging [LY13]. mesh [MYX^{+}14]. mesh-based [MYX^{+}14]. message [MAM^{+}16]. meta [SS93]. meta-programming [SS93]. metadata [PIR20, SSQJ99]. method [GRJD15, HK17, KJMJ10, LLS12, LWC^{+}12, WHC19, YA22]. methodology [BC94]. methods [Han95, MN18, MA00, VPB19]. metrices [SL99]. microarray [AB12]. microservice [CDT18, GWB20]. microwave [PSS17]. Migration [HCT^{+}13, SPPP11]. Migration-based [HCT^{+}13]. Migrations [OSG23, GSKJ18]. Minimization [AKBJ14]. minimize [CRX18]. minimum [BP18]. mining [FRS16, VHS20]. mirroring [MH11]. mitigating [TCPC17]. mitigation [SSE18]. Mixture [BBSB24]. MLC [HCC^{+}17, LCC^{+}21]. MLCLSP [TAdOD13]. Mobile [LMLAK17, AJS^{+}13, MAM^{+}16, VF01, WW01, ZCG^{+}16]. Model [BEL24, BBSB24, MCMS22, GDJ96, HEK^{+}13, MBR17, PHJJ15, THW14]. model-based [HEK^{+}13]. model-driven [MBR17]. Modeling [ARS15, AAR^{+}24, GLPT12, BSV97, GRJD15, KKEK17, LS21, MR01, SP16, WLK^{+}21, WA94, YAG^{+}15]. modelling [DOP^{+}14, SB16, vD02]. models [DF97, KR13, MPSR00, SPB01, WJdCS22]. Molecular [DQG^{+}23]. Monetization [BZT23]. Monitoring [BSVV96, CO00, NVP21]. mouse [LNLC16]. movie [HK17]. moving [CMDM20, MMC12]. MPC [Ish96]. Multi [YLJ24, DCUA20, FD15, HF19, JCP17, KP18, KKEK17, LW12, MZAS14, PKVM21, SWH^{+}13]. Multi-Agent [YLJ24]. multi-consumer [FD15]. multi-controller [LW12]. multi-core [SWH^{+}13]. multi-pedestrian [PKVM21]. multi-producer [FD15]. multi-robot [DCUA20]. multi-scale [KKEK17]. multi-sensor [MZAS14]. multi-step [HF19]. multi-tenant [KP18]. multi-threaded [JCP17]. multicast [PPM18]. multicore [SM22]. multimedia [BSV^{+}00, MVOdGCP11]. Multimodal [LNLC16, PKVM21, AJS^{+}13]. Multiple [SW16, AC20, Rai99]. multiprocessor [KL15, KL17, QM15].

named [YA22]. NAND [KHC^{+}11]. natural [GN22]. navigation [SSS^{+}11]. Navy [WCR95]. near [WZL^{+}22]. negotiation [FS02]. neophytes [ABLRC21]. net [XDS22]. Network [HW24, AJS^{+}13, BSV^{+}00, BSG16, CRX18, SD12, SW16, SWS20, TBC94, XDS22]. networked [Ang95]. networking [CRX18]. networks [BDWP20, CCSK11, MBR13, MYX^{+}14, RHA11, STP^{+}12, THW14, TFSO14, WLK^{+}21]. Neural [AAR^{+}24, HW24, SD12, WLK^{+}21, XDS22]. NMF [Alo18]. node [TFSO14]. NODEs [AAR^{+}24]. Non [PVT24, Bad96, KL15, MN18, SER15]. Non-Binary [PVT24]. non-coding [MN18]. non-cross-functional [SER15]. non-DVS [KL15]. non-uniform [Bad96]. note [Vil99]. novel [AHH12, APPV21, LWC^{+}12, LKP12, QHW^{+}13]. number [Baa19]. numerical [Ber95b].

O [BBSB24]. Object [SN96, CW19, MA00, NDS96]. object-oriented [MA00]. objects [MMC12, SUV97, TJ12]. observations [CMDM20]. observer [CCSK11]. obstacle [SSS^{+}11]. occupant [LMLAK17]. offs [KP18]. on-line [PND^{+}13]. One [DQG^{+}23, FD12, KWBE20]. One-scan [FD12]. One-Shot [DQG^{+}23]. onine [BSG16]. online [SD13]. Ontology [Baa23, CP14, KP18, FdE11, RFB^{+}15]. Ontology-based [KP18]. OO [MZ97]. OODBMSs [WA94]. OOlong [CW19]. OOP [Rin00]. Open [WCS20]. Open-source [WCS20]. Operations [Baa23, LV18]. operative [KK97]. operator [VFMM94]. operators [MVOdGCP11]. Opinions [BVC^{+}14]. Opportunities [CYLL16]. Optimal [Baa23, GSKJ18]. optimization [AdSR^{+}12, MPNdL10, PKBH17, SER15, SSQJ99]. Optimizing [DOP^{+}14, LLC16]. Oracle [MMC12]. Order [YLJ24]. orders [Ros98, Ros19]. Ordinary [AAR^{+}24]. organization [LDY20]. Oriented [SN96, ARS15, LCM18, LCL12, MA00, MVOdGCP11, NDS96]. overflow [LLS12, SHV13]. overtime [TAdOD13]. Overview [SN96, Jan93]. Oxidase [BP18].

P2P [THW14]. packets [CCSK11]. packing [Rai99]. page [SPPP11]. papers [Dam01b]. Parallel [SN96, Bad96, Chi96, GG11, Ish96, Kes96, KL17, LLMW93, LV18, NDS96]. parallelizing [ATS96]. parameterized [CP14]. parsing [Suc99]. participation [MAM^{+}16]. partition [ATS96]. partitioning [AC20]. party [AHH12]. passive [THW14]. past [WCS20]. patching [SHV13]. path [LWK14, PKVM21]. pattern [FRS16, RFB^{+}15]. patterns [CP14, IZ21, Val97]. PCM [HCC^{+}17]. pedestrian [PKVM21]. people [AGH^{+}21, BVC^{+}14]. perceived [MCY^{+}19]. performance [Chi96, GWB20, HURU11, JPM16, Kes96, LW12, OD93, WLK^{+}21]. peripheral [VWT13]. personal [NYKI98]. personalizing [KR13]. persuasive [PHJJ15]. phosphorylation [MBP12]. physical [BDWP20, LCC^{+}21, WA94]. picture [PIR20]. pipeline [KTH18]. pipeline-based [KTH18]. platform [HMY95, KWBE20, WLK^{+}21]. platform-aware [WLK^{+}21]. platforms [LCM18, SWH^{+}13]. Playing [SKEC16]. pointcloud [XDS22]. Polarization [LOCD23]. policies [PEVR13]. Pose [UBA^{+}23]. positioning [YKdSM14]. Post [HK15]. Postmortem [HAG21]. Power [AAR^{+}24, VWT13]. practice [SAL19]. Practices [PVT24, Ben00]. pragmatic [HMY95]. pre [DL16]. pre-compilation [DL16]. Prediction [AGH^{+}21, AHH12, APPV21, CDAR11, MZAS14, MBP12, PKVM21, SPPP11]. preference [NL21]. preference-based [NL21]. prefetching [HCC^{+}15]. preprocessed [SWS20]. preprocessing [CMDM20]. preprocessor [SW99]. present [Raj99b, WCS20]. Preserving [MCMS22, RHA11]. prevention [FS12, MZAS14]. principle [LDY20]. prior [BCI18]. Prioritization [LSPC23]. priority [OD93]. PRIUS [LSPC23]. privacy [AHH12, BVC^{+}14, PZ18, RHA11, SGPM22]. private [LK13]. Probabilistic [MBP12]. problem [LK93, LV18, Rai99, SER15, TFSO14]. process [AM00, BSV97, Ben00, CO00, HF19, KKH^{+}17, KR13, LLC16, PC19, RFB^{+}15]. processes [FS02]. processing [Jan93, KK97, MBR13, OKK13, STP^{+}12, VFMM94, WW01, WCR95]. processors [Bad96]. producer [FD15]. product [NL21]. products [GN22, KMR20]. profiles [BSG16]. program [BSVV96, Ros98, SS02]. programmable [BC94]. Programming [SN96, BDMT16, FS12, FS02, NDS96, SS93, SKH93, VHS20]. programs [HAG21, SHV13, YSCX17]. Projects [SAG23]. Prolog [Cit93, Ros98, SS93]. Prone [OSG23]. Protein [CDAR11, MBP12]. protocol [CJ10, OD93, SW16]. protocols [MYX^{+}14]. provisioning [GSKJ18]. proxy [MH11]. pseudo [VPB19]. pseudo-relevance [VPB19]. publication [HMY95]. publish [UOB^{+}13]. publish-subscribe [UOB^{+}13]. pull [PPM18]. push [PPM18, YHTN01]. push-based [YHTN01].

quadratic [Hua93]. qualitative [Ber95b]. Quality [DFP^{+}22, AGH^{+}21, BWLP01, BSG16, HDC^{+}14, KP18]. quality-of-service [HDC^{+}14]. Quantile [YH16]. queries [FT19, GN22, GG11, MBR13]. query [STP^{+}12, SSQJ99]. Querying [MPFC12, DL16]. Question [ÇÖT24]. Quick [CGR99]. Quick-tests [CGR99].

Racket [PKBH17]. RACS [BZT23]. radio [ZPH^{+}15]. RAM [LCC^{+}21]. random [JPM16]. Range [ATS96, STP^{+}12]. ranking [HK15, RT20]. rapid [AM00, CPT10]. ratings [AF21, LWC^{+}12]. RDF [MPFC12]. Re [Dev99]. Re-targetability [Dev99]. reading [LNLC16]. Real [ABCW95, SAL19, UBA^{+}23, AC20, BSV^{+}00, CMDM20, FS12, KL17, LWK14, OTYM01, QM15, QHW^{+}13, SM22, TBC94, WZL^{+}22, XDS22]. Real-time [SAL19, AC20, BSV^{+}00, FS12, KL17, LWK14, QM15, QHW^{+}13, SM22, WZL^{+}22, XDS22]. real-world [CMDM20]. realization [HF19]. reasoned [BV99]. reasoning [BCI18, QHW^{+}13]. Recognition [PVT24, HURU11, YA22]. recommendation [PHJJ15, TK22]. recommender [AF21]. reconfigurable [EHBS21]. Record [PKBH17, UOB^{+}13]. recovery [FdE11, GD14, HAG21, vD02]. redistricting [LK93]. reducing [KJMJ10]. redundancy [LY13]. Refaster [OSG23]. reflective [CCDS13]. Regex [BDMT16]. Regex-based [BDMT16]. regions [CMDM20]. Register [KJMJ10]. Register-relocation [KJMJ10]. regression [MPNdL10]. regular [TJ12]. rehabilitation [DGS^{+}14]. reification [GDJ96]. Reinforcement [DQG^{+}23, BCI18]. related [APSHB15, WdCS21]. Relating [Baa23]. relational [YLY^{+}22]. relations [SL99]. relations-based [SL99]. relevance [MBP12, VPB19]. relevant [AB12]. reliability [Ano01b]. reliable [RHA11]. relocation [KJMJ10]. remotely [Lea98]. removal [ZSRS13]. renaming [KJMJ10]. rendering [KTH18, TH19]. renewed [Suc99]. Repairs [Baa23]. report [vD02]. repositories [SSQJ99]. representation [Ber95b, MN18]. Reputation [BZT23, MPNdL10]. Reputation-Aware [BZT23]. requirement [FS02]. requirements [FRNNS18, VGH^{+}16]. resident [LF19]. resilient [KRHH21]. resolution [NHM15]. resource [HXGB22, LCL12]. resource-bounded [LCL12]. response [DGS^{+}14, LLC16]. RESTful [RP11]. restrictions [Baa19]. resulting [DAI^{+}21]. resume [LLC16]. resuming [HCC^{+}15]. retrieval [ABLRC21, AdSR^{+}12, BS97, SUV97]. retrospective [VGH^{+}16]. return [SB96]. reusability [CBV97]. reusable [BS97, CP14, SUV97]. Reuse [BSV^{+}00, KGKK96, BSVV96, CW19, DF96, DF97, KK97, SB96, Val97]. reused [GS96]. reverse [SSW02, SS02]. review [AF21, KMR20]. reviews [SD13]. revisited [KWBE20]. ring [FD15]. risk [MZAS14]. RMI [RM12]. RNA [MN18, SHP22]. road [AGH^{+}21]. robot [DCUA20]. Robust [HURU11, LK13, MH11]. robustness [HEK^{+}13]. Role [BEL24]. ROM/WWW [HMY95]. Rot [LFT23]. rotation [ABCW95]. routing [CCSK11, HDC^{+}14]. RSA [SWS20]. RSSI [YKdSM14]. RSSI-based [YKdSM14]. RTRS [QHW^{+}13]. rule [FD12, SHV13]. rule-based [SHV13]. rules [AB12, QHW^{+}13]. runtime [GD14].

SAC [AAR^{+}24, Dam01b, VGH^{+}16]. safety [FMNP94, Gea94, NRH13]. safety-critical [NRH13]. SAGA [XDS22]. SAGA-net [XDS22]. sampling [PND^{+}13]. SART [STP^{+}12]. Scalability [Det01]. Scale [SAG23, Alo18, JCP17, KKEK17, LBKV01, RC99, TCPC17]. scaling [ABCW95]. scan [FD12]. ScanMe [ZCG^{+}16]. scenario [SSE18]. scenarios [HURU11]. schedule [TCPC17]. scheduler [SM22, SWH^{+}13]. Scheduling [KL17, YHTN01, AC20, LV18, QM15]. scheme [CCSK11, HCC^{+}15, LKP12, LWK14, TFSO14]. Science [Dij99]. Scientific [SN96, Bad96, Vil99, WdCS21, WJdCS22]. scoring [BP18]. Scripting [JL99]. Scrum [SER15]. Second [YLJ24, FS02]. Second-Order [YLJ24]. secondary [CDAR11]. secure [LK13, OWB^{+}14, YKdSM14]. Security [Ang95, LU95, Han95, VF01]. Seed [SH23]. segments [BP18]. Selecting [GSKJ18, BBH11]. Selection [BZT23, TD14, YPB19]. Selector [LFT23]. self [ARS15, GD14]. self-adaptive [ARS15]. self-recovery [GD14]. Semantic [GAB^{+}23, BSG16, FS12, GG11, NOVS11, SSR21, UOB^{+}13, WdCS21]. semantic-based [SSR21]. semi [CGP02]. semi-structured [CGP02]. sensed [Lea98]. sensing [ZPH^{+}15]. sensor [CCSK11, MBR13, MZAS14, RHA11, STP^{+}12, TFSO14]. Sentiment [LFT23, GN22, SD12, SD13]. sequence [MN18, SHP22]. sequences [BP18, Hua93]. sequential [APPV21]. Series [PeSNY23]. server [MH11]. servers [AW95]. Service [LCL12, ARS15, BBH11, HDC^{+}14, RP11, ZCG^{+}16]. Service-oriented [LCL12, ARS15]. services [CRX18]. Session [Dam01b, Raj99c]. sets [Alo18, PPM18]. setup [KMR20]. Shade [CR98]. shape [XDS22]. shape-assisted [XDS22]. sharding [KRHH21]. shoaling [Lea98]. shop [LV18]. Shot [DQG^{+}23, YLY^{+}22]. side [LF19, TO94]. side-channel [LF19]. side-stick [TO94]. Signature [SMK17]. significant [FD12]. similarity [HK17]. SimPal [NRH13]. simplifying [SL99]. Simulating [Lea98]. Simulation [Cit93, QM15, KGKK96]. Simulation-based [QM15]. simulations [Ber95b]. simulators [CJ10]. sites [MBP12, Van02]. situ [AJS^{+}13]. Situations [DQG^{+}23]. Size [SWS20]. Sketching [PeSNY23]. SLC [HCC^{+}17]. SLC-MLC [HCC^{+}17]. slices [YSCX17]. SLYMFAST [BAK22]. Smart [BEL24, LMLAK17, MYX^{+}14, YAG^{+}15]. SmartStream [KRHH21]. smells [DAI^{+}21]. SMT [FRNNS18]. SMT-based [FRNNS18]. social [AJS^{+}13, BSG16, HK17, PHJJ15]. Soft [AAR^{+}24]. Software [Bad96, ÇÖT24, MPSR00, SAG23, vD02, BS97, BWLP01, CRX18, CO00, Dev99, FS02, GS96, KKH^{+}17, KK97, NL21, NRH13, OWB^{+}14, RFB^{+}15, VSF97, Val97, WW01]. software-defined [CRX18]. solid [LW12]. solid-state [LW12]. Solutions [PVT24]. solve [SER15]. solving [AM98]. source [WCS20]. space [AdSR^{+}12, Hua93, LLC16, MPFC12]. SPARQL [FT19, GG11]. sparse [Alo18, VFMM94]. spatiotemporal [MMC12]. Special [Dam01b, Raj99c, SN96, Val97]. specific [PZBdBC18, SB16]. specification [EHBS21]. specifications [Cit93]. spectrum [ZPH^{+}15]. speculation [KRHH21]. speeding [STP^{+}12]. spread [APPV21]. sprint [SER15]. SQL [DL16]. stability [MRS12]. stack [CJ10]. stacked [CYLL16]. standard [Gea94]. standards [Bou95, LSV12, RFB^{+}15]. State [AAR^{+}24, HEK^{+}13, HAG21, LW12]. stationary [YH16]. statistical [Han95, YAG^{+}15]. step [HF19]. stepwise [MPNdL10]. stick [TO94]. STL [ATS96, KCDS98]. stochastic [YH16]. storage [HCT^{+}13, PND^{+}13, WHC19]. Stories [LSPC23]. storm [MBR17]. storm-based [MBR17]. strategies [CMDM20, DGS^{+}14, YHTN01]. strategy [LC13]. stream [OKK13]. streaming [KRHH21, MBR17]. streams [YPB19]. stroke [SMK17]. structure [MN18, STP^{+}12]. structured [CGP02, Tre02]. structures [CDAR11, PKBH17]. STT [LCC^{+}21]. student [MAM^{+}16, MCY^{+}19]. Study [LWS^{+}23, AW95, BC94, DAI^{+}21, IZ21, KMR20, MAM^{+}16, MRS12, NRH13, PZBdBC18, RZW01, VGM^{+}20]. subscribe [UOB^{+}13]. subset [TD14, YPB19]. substations [VFMM94]. subsumption [LY13]. suitable [GSKJ18]. summaries [GN22]. Summary [TJ12]. Summary-based [TJ12]. Super [NHM15]. supervision [HMB12]. support [AGH^{+}21, SB16, Tre02, WCR95]. Supporting [FS12, HYL^{+}17, Van02]. SVM [SD13]. swap [LLC16]. swarm [YKdSM14]. Swedish [SAG23]. Sybil [SKEC16]. symmetric [SWS20]. synergistically [CYLL16]. synopsis [BV99]. synthetic [CGP02, KTH18, TH19]. System [AAR^{+}24, LWS^{+}23, UBA^{+}23, BDWP20, BC94, JCP17, LKH17, LLC16, QHW^{+}13, RZW01, SP16, SB16, SKH93, TBC94, VFMM94, Van02, WCR95, WJdCS22]. systematic [DAI^{+}21]. Systems [ÇÖT24, YLJ24, Ano01b, AF21, BHK^{+}15, CC17, CGR99, EHBS21, HCC^{+}15, HCC^{+}17, HCT^{+}13, KKEK17, KGKK96, KL15, KL17, LK93, LCC^{+}21, LCL12, LKP12, MPNdL10, NOVS11, NVP21, PEVR13, QM15, RC99, SSR21, SW16, SGPM22, VF01, WHC19, YHTN01].

tag [LWC^{+}12]. tale [DCUA20]. Taming [BAK22]. targetability [Dev99]. Task [KL15, DCUA20]. tasks [AC20, KL17]. TCP [LU95, TCPC17]. TCP/IP [LU95]. TD [ÇÖT24]. teaching [LDY20]. teams [SER15]. Technical [ÇÖT24, DAI^{+}21]. technique [Alo18]. techniques [TD14]. Technologies [GAB^{+}23]. technology [Ano01b, HMY95]. telecommunications [Nyg94]. telephony [TBC94]. temperature [KJMJ10, SM22]. Template [SSS^{+}11]. Template-based [SSS^{+}11]. temporal [APSHB15]. tenant [KP18]. term [VPB19]. term-based [VPB19]. termination [WCR95]. testing [BK19, GD14, HEK^{+}13]. tests [CGR99]. text [AF21, RT20]. theoretical [KKH^{+}17]. theory [MKR93, SAL19, YH16]. thermal [CYLL16, KJMJ10]. thermal-aware [KJMJ10]. Things [LWK14]. third [AHH12]. threaded [JCP17]. Threatening [GAB^{+}23]. three [Hua93]. Tigrinya [YA22]. Time [PeSNY23, UBA^{+}23, YLJ24, AC20, ABCW95, BSV^{+}00, FS12, FF11, KL17, LWK14, QM15, QHW^{+}13, SAL19, SM22, TBC94, WZL^{+}22, XDS22]. Time-Domain [PeSNY23]. Time-Series [PeSNY23]. tolerance [RM12]. tomography [PSS17]. tool [CPT10, SB16]. tools [Dev99, SSW02, Suc99, WCS20]. top [BBH11]. top- [BBH11]. tour [SKH93]. track [VGH^{+}16]. trade [KP18]. trade-offs [KP18]. traffic [LCL12]. Training [BBSB24, HW24]. Trans [PVT24]. Trans-Inclusive [PVT24]. transactional [HLF^{+}19]. transfer [MZ97]. transform [FF11].

Transforming [Rin00]. transient [WZL^{+}22]. translating [MH11]. translation [HLF^{+}19, TWK^{+}13]. translators [BV99]. Transparently [RM12]. transport [SW16]. traversal [FT19]. treatment [LBKV01, SGPM22]. tree [STP^{+}12]. trends [WCS20]. triad [MKR93]. Triggered [YLJ24]. triploid [LHO12]. Trust [TK22, BVC^{+}14]. Trust-aware [TK22]. trusted [AHH12]. TuCSoN [NOVS11]. tuned [WJdCS22]. tunnel [LK13]. tuple [NOVS11]. Two [SAG23, DCUA20, WA94]. types [RT20].

UAV [YKdSM14]. UK [TBC94]. UML [AM00]. uncertain [FdE11]. Understanding [KMR20, KKH^{+}17, CDT18, SS02]. understands [TJ12]. unified [Ber95b, HMB12]. uniform [Bad96]. units [MBP12]. unregulated [SPB01]. unsolicited [TD14]. upon [LSV12]. urban [LCL12, SSS^{+}11]. Usability [PZBdBC18, KMR20]. usage [PKBH17, YAG^{+}15]. use [SPD^{+}19, Vil99]. User [LSPC23, MCMS22, BK19, CCDS13, GN22, LWC^{+}12, MR01, TK22]. user-based [LWC^{+}12]. Using [MAM^{+}16, MMSV02, MCMS22, MA22, PEVR13, SS93, SH23, AJS^{+}13, AAR^{+}24, BP18, CCSK11, CO00, Cit93, DL16, HK17, JCP17, MN18, MPNdL10, MA00, MBP12, OTYM01, PND^{+}13, PPM18, RHA11, SD12, SKH93, VHS20, VGM^{+}20, WJdCS22, WHC19, YH16, YKdSM14, YPB19]. USRP [ZPH^{+}15].

validating [ARS15]. validation [EHBS21]. value [FD12, YSCX17]. value-flow [YSCX17]. values [LY13]. variability [BDWP20, LCM18]. variability-intensive [LCM18]. various [SB16]. vehicle [FD12]. verification [MBR17, SMK17]. Versions [Bou95]. via [DOP^{+}14, HLF^{+}19, LF19, MVOdGCP11, YLJ24]. video [DGS^{+}14, MCY^{+}19]. view [SUV97]. viewer [NYKI98]. views [CGP02]. Villasimius [FS02]. Virtual [CJ10, GSKJ18, LF19, OTYM01, SWH^{+}13, VWT13]. visual [LKH17]. visualization [LBKV01]. visualizing [NYKI98]. vital [APSHB15]. voice [NL21]. voltage [CYLL16]. voltage-frequency [CYLL16]. VR [KMR20]. vs [AC20]. Vulnerability [BEL24].

W3C [LSV12]. wait [FD15]. wait-free [FD15]. watermarking [HURU11]. waves [Lea98]. WCRE [vD02]. web [FT19, GG11, LSV12, LS21, MH11, MCP15, RP11, Van02, AW95, BBH11, Ber95a, Men95, War95]. web-based [MCP15]. webmaster [War95]. weighted [Rai99]. where [IZ21]. whole [MH11]. Wide [LWS^{+}23, War95]. wire [TO94]. wireless [CCSK11, LKP12, MBR13, RHA11, SWS20, TFSO14]. without [AHH12]. Word [LFT23]. workload [AW95, MCY^{+}19]. workshop [FS02]. workspaces [BHK^{+}15]. world [CMDM20, War95]. worlds [OTYM01]. worm [THW14]. Write [HCC^{+}17, JPM16]. Write-aware [HCC^{+}17]. WS [CPT10]. WS-BPEL [CPT10]. WSN [AKBJ14, CJ10]. WWW [HMY95].

XFormsDB [LSV12]. XML [CGP02, LC13, MMSV02, Nic02, Tre02, Van02]. XP [FS02]. XP2001 [FS02].

yield [WCR95].

zero [FT19]. zero-knowledge [FT19].

References

Aslami:2024:PSF

[AAR^{+}24] Pooja Aslami, Tara Aryal, Astha Rai, Niranjan Bhujel, Hossein Moradi Rekabdarkolaee, Kaiqun Fu, Reinaldo Tonkoski, Zongjie Wang, and Timothy M. Hansen. Power system frequency dynamics modeling, state estimation, and control using neural ordinary differential equations (NODEs) and soft actor– critic (SAC) machine learning approaches. ACM SIGAPP Applied Computing Review, 24(1):24–39, March 2024. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/10.1145/3663652.3663655.

Antonie:2012:BRA

[AB12] Luiza Antonie and Kyrylo Bessonov. Biologically relevant association rules for classification of microarray data. ACM SIGAPP Applied Computing Review, 12(1):12–23, April 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2188379.2188381.

Andrews:1995:RTC

[ABCW95] David L. Andrews, Randy Brown, Charles Caldwell, and Andrew Wheeler. Real time character scaling and rotation. ACM SIGAPP Applied Computing Review, 3(1):19–22, June 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
214310.214434.

Audeh:2021:IEI

[ABLRC21] Bissan Audeh, Michel Beigbeder, Christine Largeron, and Diana Ramírez-Cifuentes. Improving exploratory information retrieval for neophytes: machine learning approach with feature analysis. ACM SIGAPP Applied Computing Review, 20(4):50– 64, January 2021. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3447332.3447336.

Abeni:2020:ESR

[AC20] Luca Abeni and Tommaso Cucinotta. EDF scheduling of realtime tasks on multiple cores: adaptive partitioning vs. global scheduling. ACM SIGAPP Applied Computing Review, 20(2): 5–18, July 2020. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3412816.3412817.

Avalhais:2012:FSO

[AdSR^{+}12] Letricia P. S. Avalhais, Sergio F. da Silva, Jose F. Rodrigues, Agma J. M. Traina, and Caetano Traina. Feature space optimization for content-based image retrieval. ACM SIGAPP Applied Computing Review, 12(3):7–19, September 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2387358.2387359.

Aslanyan:2021:LLJ

[AF21] Tatev Karen Aslanyan and Flavius Frasincar. LDA-LFM: a joint exploitation of review text and ratings in recommender systems. ACM SIGAPP Applied Computing Review, 21(2):33– 47, June 2021. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3477127.3477130.

Antoniadi:2021:PQL

[AGH^{+}21] Anna Markella Antoniadi, Miriam Galvin, Mark Heverin, Orla Hardiman, and Catherine Mooney. Prediction of quality of life in people with ALS: on the road towards explainable clinical decision support. ACM SIGAPP Applied Computing Review, 21(2):5–17, June 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3477127.3477128.

Ahamed:2012:NLP

[AHH12] Sheikh Iqbal Ahamed, Md. Munirul Haque, and Chowdhury Sharif Hasan. A novel location privacy framework without trusted third party based on location anonymity prediction. ACM SIGAPP Applied Computing Review, 12(1):24–34, April 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2188379.2188382.

Adibuzzaman:2013:SAD

[AJS^{+}13] Mohammad Adibuzzaman, Niharika Jain, Nicholas Steinhafel, Munir Haque, Ferdaus Ahmed, Sheikh Ahamed, and Richard Love. In situ affect detection in mobile devices: a multimodal approach for advertisement using social network. ACM SIGAPP Applied Computing Review, 13(4):67–77, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2577554.2577562.

Ahmed:2014:ECB

[AKBJ14] Syed Hassan Ahmed, Dongkyun Kim, Safdar Hussain Bouk, and Nadeem Javaid. Error control based energy minimization for cooperative communication in WSN. ACM SIGAPP Applied Computing Review, 14(3):55–64, September 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2670967.2901789.

Alostad:2018:NBS

[Alo18] Jasem M. Alostad. NMF based sparse Cholesky decomposition technique for dimensionality scale back in large data sets. ACM SIGAPP Applied Computing Review, 18(1):41–49, April 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3212069.3212073.

Arbab:1998:CHD

[AM98] Farhad Arbab and Eric Monfroy. Coordination of heterogeneous distributed cooperative constraint solving. ACM SIGAPP Applied Computing Review, 6(2):4–17, September 1998. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
297114.297115.

Armano:2000:RDP

[AM00] Giuliano Armano and Michele Marchesi. A rapid development process with UML. ACM SIGAPP Applied Computing Review, 8(1):4–11, September 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/361651.361653.

Angaye:1995:SNE

[Ang95] Cleopas O. Angaye. Security in a networked environment. ACM SIGAPP Applied Computing Review, 3(1):2–5, June 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
214310.214312.

Anonymous:2001:IE

[Ano01a] Anonymous. Improving engineering. ACM SIGAPP Applied Computing Review, 9(2):13–16, July 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/512000.512003.

Anonymous:2001:IRC

[Ano01b] Anonymous. Increasing the reliability of control systems with agent technology. ACM SIGAPP Applied Computing Review, 9(2):6–12, July 2001. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/512000.512002.

Aragona:2021:LSP

[APPV21] Dario Aragona, Luca Podo, Bardh Prenkaj, and Paola Velardi. Latent and sequential prediction of the novel coronavirus epidemiological spread. ACM SIGAPP Applied Computing Review, 21(3):5–18, September 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/10.1145/3493499.3493500.

Abbes:2015:WTE

[APSHB15] Rafik Abbes, Karen Pinel-Sauvagnat, Nathalie Hernandez, and Mohand Boughanem. When temporal expressions help to detect vital documents related to an entity. ACM SIGAPP Applied Computing Review, 15(3):49–58, October 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2835260.2835263.

Arcaini:2015:MVS

[ARS15] Paolo Arcaini, Elvinia Riccobene, and Patrizia Scandurra. Modeling and validating self-adaptive service-oriented applications. ACM SIGAPP Applied Computing Review, 15(3):35– 48, October 2015. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2835260.2835262.

Austern:1996:RPA

[ATS96] Matthew H. Austern, Ross A. Towle, and Alexander A. Stepanov. Range partition adaptors: a mechanism for parallelizing STL. ACM SIGAPP Applied Computing Review, 4 (1):5–6, April 1996. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/240732.240734.

Arlitt:1995:WCS

[AW95] Martin F. Arlitt and Carey L. Williamson. A workload characterization study of Internet Web servers. ACM SIGAPP Applied Computing Review, 3(2):1–4, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/228228.228229.

Baader:2019:ECR

[Baa19] Franz Baader. Expressive cardinality restrictions on concepts in a description logic with expressive number restrictions. ACM SIGAPP Applied Computing Review, 19(3):5–17, November 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3372001.3372002.

Baader:2023:ROR

[Baa23] Franz Baader. Relating optimal repairs in ontology engineering with contraction operations in belief change. ACM SIGAPP Applied Computing Review, 23(3):5–18, September 2023. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3626307.3626308.

Baden:1996:SIN

[Bad96] Scott B. Baden. Software infrastructure for non-uniform scientific computations on parallel processors. ACM SIGAPP Applied Computing Review, 4(1):7–10, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/240732.240736.

Belcaid:2022:TDC

[BAK22] Mahdi Belcaid, Cedric Arisdakessian, and Yuliia Kravchenko. Taming DNA clustering in massive datasets with SLYMFAST. ACM SIGAPP Applied Computing Review, 22(1):15–23, March 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3530043.
3530045.

Benouaret:2011:FFS

[BBH11] Karim Benouaret, Djamal Benslimane, and Allel Hadjali. A fuzzy framework for selecting top-k Web service compositions. ACM SIGAPP Applied Computing Review, 11(3):32–40, August 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2034594.2034597.

Bouzouad:2024:AGM

[BBSB24] Meriem Bouzouad, Yasmine Benhamadi, Camélia Slimani, and Jalil Boukhobza. Adapting Gaussian mixture model training to embedded/edge devices: a low I/O, deadline-aware and energy efficient design. ACM SIGAPP Applied Computing Review, 24(2):5–18, June 2024. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3687251.3687252.

Broomfield:1994:HIP

[BC94] E. J. Broomfield and P. W. H. Chung. Hazard identification in programmable system: a methodology and case study. ACM SIGAPP Applied Computing Review, 2(1):7–14, March 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
381766.381768.

Bougie:2018:CDR

[BCI18] Nicolas Bougie, Li Kai Cheng, and Ryutaro Ichise. Combining deep reinforcement learning with prior knowledge and reasoning. ACM SIGAPP Applied Computing Review, 18(2):33–45, July 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3243064.3243067.

Bartoli:2016:RBE

[BDMT16] Alberto Bartoli, Andrea De Lorenzo, Eric Medvet, and Fabiano Tarlao. Regex-based entity extraction with active learning and genetic programming. ACM SIGAPP Applied Computing Review, 16(2):7–15, August 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2993231.2993232.

Brings:2020:AGV

[BDWP20] Jennifer Brings, Marian Daun, Thorsten Weyer, and Klaus Pohl. Analyzing goal variability in cyber-physical system networks. ACM SIGAPP Applied Computing Review, 20(2):19– 35, July 2020. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3412816.3412818.

Boi:2024:SCV

[BEL24] Biagio Boi, Christian Esposito, and Sokjoon Lee. Smart contract vulnerability detection: The role of large language model (LLM). ACM SIGAPP Applied Computing Review, 24(2):19– 29, June 2024. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3687251.3687253.

Benedicenti:2000:IPP

[Ben00] Luigi Benedicenti. Industrial practices for process improvement. ACM SIGAPP Applied Computing Review, 8(1):2–3, September 2000. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/361651.361652.

Berghel:1994:EC

[Ber94] Hal Berghel. From the Editor-in-Chief. ACM SIGAPP Applied Computing Review, 2(2):1, September 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/381777.382058.

Berghel:1995:IDW

[Ber95a] Hal Berghel. The inevitable demise of the Web. ACM SIGAPP Applied Computing Review, 3(2):5–8, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/228228.228231.

Berleant:1995:URN

[Ber95b] Daniel Berleant. A unified representation for numerical and qualitative simulations. ACM SIGAPP Applied Computing Review, 3(1):23–26, June 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/214310.214435.

Badouel:2015:AWD

[BHK^{+}15] Eric Badouel, Loïc Hélouët, Georges-Edouard Kouamou, Christophe Morvan, and Nsaibirni Robert Fondze. Active workspaces: distributed collaborative systems based on guarded attribute grammars. ACM SIGAPP Applied Computing Review, 15(3):6–34, October 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2835260.2835261.

Bunder:2019:TBD

[BK19] Hendrik Bünder and Herbert Kuchen. Towards behavior-driven graphical user interface testing. ACM SIGAPP Applied Computing Review, 19(2):5–17, August 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3357385.3357386.

Bouvier:1995:VSH

[Bou95] Dennis J. Bouvier. Versions and standards of HTML. ACM SIGAPP Applied Computing Review, 3(2):9–15, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
228228.228232.

Belcaid:2018:DAC

[BP18] Mahdi Belcaid and Guylaine Poisson. Detecting anomalies in the Cytochrome C Oxidase I amplicon sequences using minimum scoring segments. ACM SIGAPP Applied Computing Review, 17(4):6–14, January 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3183628.3183629.

Baruchelli:1997:FAF

[BS97] Francesco Baruchelli and Giancarlo Succi. A fuzzy approach to faceted classification and retrieval of reusable software components. ACM SIGAPP Applied Computing Review, 5(1):15–20, June 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
270895.270897.

Besel:2016:QSI

[BSG16] Christoph Besel, Jörg Schlötterer, and Michael Granitzer. On the quality of semantic interest profiles for onine social network consumers. ACM SIGAPP Applied Computing Review, 16(3):

5–14, November 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3015297.3015298.

Benedicenti:1997:PMD

[BSV97] Luigi Benedicenti, Giancarlo Succi, and Tullio Vernazza. From process modeling to domain modeling. ACM SIGAPP Applied Computing Review, 5(2):28–32, September 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/297075.297087.

Benedicenti:2000:RLR

[BSV^{+}00] Luigi Benedicenti, Giancarlo Succi, Tullio Vernazza, George Kovacs, and Andrea Valerio. Reuse libraries for real-time multimedia over the network. ACM SIGAPP Applied Computing Review, 8(1):12–19, September 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/361651.361654.

Benedicenti:1996:MER

[BSVV96] L. Benedicenti, G. Succi, A. Valerio, and T. Vernazza. Monitoring the efficiency of a reuse program. ACM SIGAPP Applied Computing Review, 4(2):8–14, September 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/251560.251582.

Benedicenti:1999:LTR

[BV99] Luigi Benedicenti and Tullio Vernazza. Language translators: a reasoned synopsis. ACM SIGAPP Applied Computing Review, 7(3):4–10, September 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/333630.333632.

Basu:2014:OPF

[BVC^{+}14] Anirban Basu, Jaideep Vaidya, Juan Camilo Corena, Shinsaku Kiyomoto, Stephen Marsh, Guibing Guo, Jie Zhang, and Yutaka Miyake. Opinions of people: factoring in privacy and trust. ACM SIGAPP Applied Computing Review, 14(3):7– 21, September 2014. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2670967.2670968.

Benedicenti:2001:EQC

[BWLP01] Luigi Benedicenti, Victor Wei Wang, Peter Lee, and Raman Paranjape. Establishing quality control in software agents. ACM SIGAPP Applied Computing Review, 9(3):31– 33, September 2001. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/570132.570137.

Batool:2023:BRT

[BZT23] Zahra Batool, Kaiwen Zhang, and Matthew Toews. Block-RACS: Towards reputation-aware client selection and monetization mechanism for federated learning. ACM SIGAPP Applied Computing Review, 23(3):49–65, September 2023. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3626307.3626311.

Cardino:1997:EFR

[CBV97] Guido Cardino, Francesco Baruchelli, and Andrea Valerio. The evaluation of framework reusability. ACM SIGAPP Applied Computing Review, 5(2):21–27, September 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/297075.297085.

Cemus:2017:AEB

[CC17] Karel Cemus and Tomas Cerny. Automated extraction of business documentation in enterprise information systems. ACM SIGAPP Applied Computing Review, 16(4):5–13, January 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3040575.3040576.

Cerny:2013:ADD

[CCDS13] Tomas Cerny, Karel Cemus, Michael J. Donahoo, and Eunjee Song. Aspect-driven, data-reflective and context-aware user interfaces design. ACM SIGAPP Applied Computing Review, 13(4):53–66, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2577554.2577561.

Choo:2011:EHB

[CCSK11] Hyunseung Choo, Mooshik Choi, Minhan Shon, and Doogsoo Stephen Kim. Efficient hole bypass routing scheme using observer packets for geographic routing in wireless sensor networks. ACM SIGAPP Applied Computing Review, 11(4): 7–16, December 2011. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2107756.2107757.

Chamorro:2011:PSS

[CDAR11] Alfonso E. Márquez Chamorro, Federico Divina, and Jesús S. Aguilar-Ruiz. Protein secondary structures prediction based on evolutionary computation. ACM SIGAPP Applied Computing Review, 11(4):17–25, December 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2107756.2107758.

Cerny:2018:CUM

[CDT18] Tomas Cerny, Michael J. Donahoo, and Michal Trnka. Contextual understanding of microservice architecture: current and future directions. ACM SIGAPP Applied Computing Review, 17(4):29–45, January 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3183628.3183631.

Cannataro:2002:KMX

[CGP02] Mario Cannataro, Antonella Guzzo, and Andrea Pugliese. Knowledge management and XML: derivation of synthetic views over semi-structured data. ACM SIGAPP Applied Computing Review, 10(1):33–36, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/568235.568242.

Cormier:1999:QTC

[CGR99] Rick Cormier, Ed Guy, and David E. Ruddock. Quick-tests for characterizing distributed systems. ACM SIGAPP Applied Computing Review, 7(1):5–8, April 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/570150.570153.

Chien:1996:ICD

[Chi96] Andrew A. Chien. ICC++ — a C++ dialect for high performance parallel computing. ACM SIGAPP Applied Computing Review, 4(1):19–23, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/240732.240740.

Citrin:1993:SCA

[Cit93] Wayne Citrin. Simulation of communications architecture specifications using Prolog. ACM SIGAPP Applied Computing Review, 1(1):10–17, January 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/152535.152537.

Cho:2010:VPS

[CJ10] Yookun Cho and Jinman Jung. Virtual protocol stack for WSN simulators. ACM SIGAPP Applied Computing Review, 11(1):7–13, June 2010. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/1869687.1869688.

Costa:2020:EPI

[CMDM20] Rogério Luís C. Costa, Enrico Miranda, Paulo Dias, and José Moreira. Evaluating preprocessing and interpolation strategies to create moving regions from real-world observations. ACM SIGAPP Applied Computing Review, 20(2):46–58, July 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3412816.3412820.

Cimpan:2000:DSP

[CO00] Sorana Cîmpan and Flavio Oquendo. Dealing with software process deviations using fuzzy logic based monitoring. ACM SIGAPP Applied Computing Review, 8(2):3–13, December 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
373975.373979.

Caglayan:2024:ATD

[ÇÖT24] Dilek Çaglayan and Özden Özcan-Top. Advancing technical debt management in software systems with a comprehensive TD indicator and question catalog. ACM SIGAPP Applied Computing Review, 24(2):30–54, June 2024. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3687251.3687254.

Chaturvedi:2014:ODC

[CP14] Amrita Chaturvedi and T. V. Prabhakar. Ontology driven creational design patterns as parameterized and reusable components. ACM SIGAPP Applied Computing Review, 14(1):

6–19, March 2014. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2600617.2600618.

Cesari:2010:TRD

[CPT10] Luca Cesari, Rosario Pugliese, and Francesco Tiezzi. A tool for rapid development of WS-BPEL applications. ACM SIGAPP Applied Computing Review, 11(1):27–40, June 2010. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/1869687.1869690.

Ciancarini:1998:CDA

[CR98] P. Ciancarini and D. Rossi. Coordinating distributed applets with Shade. ACM SIGAPP Applied Computing Review, 6(1): 2–12, April 1998. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/297090.297094.

Chin:2018:ASD

[CRX18] Tommy Chin, Mohamed Rahouti, and Kaiqi Xiong. Applying software-defined networking to minimize the end-to-end delay of network services. ACM SIGAPP Applied Computing Review, 18(1):30–40, April 2018. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3212069.3212072.

Castegren:2019:OCO

[CW19] Elias Castegren and Tobias Wrigstad. OOlong: a concurrent object calculus for extensibility and reuse. ACM SIGAPP Applied Computing Review, 18(4):47–60, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3307624.3307629.

Chen:2016:OSA

[CYLL16] Yi-Jung Chen, Chia-Lin Yang, Pin-Sheng Lin, and Yi-Chang Lu. Opportunities of synergistically adjusting voltagefrequency levels of cores and DRAMs in CMPs with 3d-stacked DRAMs for efficient thermal control. ACM SIGAPP Applied Computing Review, 16(1):26–35, April 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2924715.2924718.

Das:2021:TDR

[DAI^{+}21] Dipta Das, Abdullah Al Maruf, Rofiqul Islam, Noah Lambaria, Samuel Kim, Amr S. Abdelfattah, Tomas Cerny, Karel Frajtak, Miroslav Bures, and Pavel Tisnovsky. Technical debt resulting from architectural degradation and code smells: a systematic mapping study. ACM SIGAPP Applied Computing Review, 21 (4):20–36, December 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3512753.3512755.

Damiani:2001:GE

[Dam01a] Ernesto Damiani. Guest editorial. ACM SIGAPP Applied Computing Review, 9(1):1, April 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/570142.570144.

Damiani:2001:SDS

[Dam01b] Ernesto Damiani. Session details: Special issue on SAC 2001 best papers. ACM SIGAPP Applied Computing Review, 9 (1):??, April 2001. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3262410.

Dutta:2020:CCB

[DCUA20] Ayan Dutta, Emily Czarnecki, Vladimir Ufimtsev, and Asai Asaithambi. Correlation clustering-based multi-robot task allocation: a tale of two graphs. ACM SIGAPP Applied Computing Review, 19(4):5–16, January 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3381307.3381308.

Damiani:2002:GE

[DEM02] Ernesto Damiani, Valerio Elia, and Mauro Madravio. Guest editorial. ACM SIGAPP Applied Computing Review, 10(1):32, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
568235.568241.

Deters:2001:SIA

[Det01] Ralph Deters. Scalability and information agents. ACM SIGAPP Applied Computing Review, 9(3):13–20, September 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
570132.570135.

Devanbu:1999:RTS

[Dev99]  Premkumar T. Devanbu.  Re-targetability in software

tools. ACM SIGAPP Applied Computing Review, 7(3):19– 26, September 1999. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/333630.333634.

Damiani:1996:DCR

[DF96] E. Damiani and M. G. Fugini. Design and code reuse based on fuzzy classification of components. ACM SIGAPP Applied Computing Review, 4(2):26–32, September 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/251560.251585.

Devanbu:1997:EFD

[DF97] Prem Devanbu and Bill Frakes. Extracting formal domain models from exsisting code for generative reuse. ACM SIGAPP Applied Computing Review, 5(1):2–14, June 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/270895.270896.

DelBuono:2022:FEQ

[DFP^{+}22] Francesco Del Buono, Guglielmo Faggioli, Matteo Paganelli, Andrea Baraldi, Francesco Guerra, and Nicola Ferro. A framework to evaluate the quality of integrated datasets. ACM SIGAPP Applied Computing Review, 22(4):5–23, December 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3584014.
3584015.

Davison:2014:ERM

[DGS^{+}14] Richard Davison, Sara Graziadio, Kholood Shalabi, Gary Ushaw, Graham Morgan, and Janet Eyre. Early response markers from video games for rehabilitation strategies. ACM SIGAPP Applied Computing Review, 14(3):36–43, September 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2670967.2670970.

Dijkstra:1999:CSA

[Dij99] Edsger W. Dijkstra. Computing science: achievements and challenges. ACM SIGAPP Applied Computing Review, 7(2): 2–9, March 1999. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/335527.335528.

Douglas:2016:ESQ

[DL16] Graeme Douglas and Ramon Lawrence. Efficient SQL querying on embedded devices using pre-compilation. ACM SIGAPP Applied Computing Review, 16(2):42–47, August 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2993231.2993235.

Duarte:2021:IC

[DM21] José Duarte and Mark McKenney. Interpolating concavities. ACM SIGAPP Applied Computing Review, 21(3):49–60, September 2021. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3493499.3493503.

Diegues:2014:OHH

[DOP^{+}14] Nuno Diegues, Muhammet Orazov, João Paiva, Luís Rodrigues, and Paolo Romano. Optimizing hyperspace hashing via analytical modelling and adaptation. ACM SIGAPP Applied Computing Review, 14(2):23–35, June 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2656864.2656866.

Dong:2023:ADD

[DQG^{+}23] Liang (Leon) Dong, Yuchen Qian, Paulina Gonzalez, Orhan K. Öz, and Xiankai Sun. Advancing drug discovery with deep learning: Harnessing reinforcement learning and one-shot learning for molecular design in low-data situations. ACM SIGAPP Applied Computing Review, 23(1):36–48, March 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/10.1145/3594264.3594267.

El-Hokayem:2021:FSV

[EHBS21] Antoine El-Hokayem, Marius Bozga, and Joseph Sifakis. A framework for the specification and validation of dynamic reconfigurable systems. ACM SIGAPP Applied Computing Review, 21(2):18–32, June 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3477127.3477129.

Florez:2012:OSR

[FD12] Omar U. Florez and Curtis Dyreson. One-scan rule extraction to explain significant vehicle interactions with guaranteed error value. ACM SIGAPP Applied Computing Review, 12(2): 27–38, June 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2340416.2340419.

Feldman:2015:WFM

[FD15] Steven Feldman and Damian Dechev. A wait-free multiproducer multi-consumer ring buffer. ACM SIGAPP Applied Computing Review, 15(3):59–71, October 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2835260.2835264.

Fanizzi:2011:COM

[FdE11] Nicola Fanizzi, Claudia d’Amato, and Floriana Esposito. Composite ontology matching with uncertain mappings recovery. ACM SIGAPP Applied Computing Review, 11(2):17–29, March 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
1964144.1964148.

Fontugne:2011:HTB

[FF11] Romain Fontugne and Kensuke Fukuda. A Hough-transform-based anomaly detector with an adaptive time interval. ACM SIGAPP Applied Computing Review, 11(3):41–51, August 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2034594.2034598.

Fenelon:1994:TIS

[FMNP94] P. Fenelon, J. A. McDermid, M. Nicolson, and D. J. Pumfrey. Towards integrated safety analysis and design. ACM SIGAPP Applied Computing Review, 2(1):21–32, March 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/381766.381770.

Filipovikj:2018:ASB

[FRNNS18] Predrag Filipovikj, Guillermo Rodriguez-Navas, Mattias Nyberg, and Cristina Seceleanu. Automated SMT-based consistency checking of industrial critical requirements. ACM SIGAPP Applied Computing Review, 17(4):15–28, January 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3183628.3183630.

Fassetti:2016:DGP

[FRS16] Fabio Fassetti, Simona E. Rombo, and Cristina Serrao. Discriminating graph pattern mining from gene expression data. ACM SIGAPP Applied Computing Review, 16(3):26– 36, November 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3015297.3015300.

Fraser:2002:XRN

[FS02] Steven Fraser and Giancarlo Succi. XP requirement negotiation workshop: co-located with XP2001 — the second international conference on eXtreme programming and agile processes in software engineering Villasimius, Cagliari, Italy, 23rd May 2001 workshop log. ACM SIGAPP Applied Computing Review, 10(1):26–31, April 2002. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/568235.568239.

Fan:2012:SSC

[FS12] Hongfei Fan and Chengzheng Sun. Supporting semantic conflict prevention in real-time collaborative programming environments. ACM SIGAPP Applied Computing Review, 12(2): 39–52, June 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2340416.2340420.

Fafalios:2019:ASQ

[FT19] Pavlos Fafalios and Yannis Tzitzikas. Answering SPARQL queries on the web of data through zero-knowledge link traversal. ACM SIGAPP Applied Computing Review, 19(3):18–32, November 2019. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/3372001.3372003.

Guittoum:2023:LST

[GAB^{+}23] Amal Guittoum, François Aïssaoui, Sébastien Bolle, Fabienne Boyer, and Noel De Palma. Leveraging semantic technologies for collaborative inference of threatening IoT dependencies. ACM SIGAPP Applied Computing Review, 23(3):32–48, September 2023. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3626307.3626310.

Gama:2014:DAF

[GD14] Kiev Gama and Didier Donsez. Deployment and activation of faulty components at runtime for testing self-recovery mechanisms. ACM SIGAPP Applied Computing Review, 14(3):44– 54, September 2014. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2670967.2670971.

Gannon:1996:HEC

[GDJ96] Dennis Gannon, Shridar Diwan, and Elizabeth Johnson. HPC++ and the Europa call reification model. ACM SIGAPP Applied Computing Review, 4(1):11–14, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/240732.240737.

Geary:1994:SSM

[Gea94] Kevin Geary. A standard for safety management in design. ACM SIGAPP Applied Computing Review, 2(1):3–6, March 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
381766.381767.

Groppe:2011:ALS

[GG11] Jinghua Groppe and Sven Groppe. Accelerating large semantic web databases by parallel join computations of SPARQL queries. ACM SIGAPP Applied Computing Review, 11(4):60– 70, December 2011. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2107756.2107762.

Gjondrekaj:2012:MAK

[GLPT12] Edmond Gjondrekaj, Michele Loreti, Rosario Pugliese, and Francesco Tiezzi. Modeling adaptation with Klaim. ACM SIGAPP Applied Computing Review, 12(4):21–35, December 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2432546.2432548.

Gao:2022:GES

[GN22] Siqi Gao and Yiu-Kai Ng. Generating extractive sentiment summaries for natural language user queries on products. ACM SIGAPP Applied Computing Review, 22(2):5–20, June 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3558053.3558054.

Gassara:2015:FMM

[GRJD15] Amal Gassara, Ismael Bouassida Rodriguez, Mohamed Jmaiel, and Khalil Drira. A formal method for modeling deployment architectures based on bigraphs. ACM SIGAPP Applied Computing Review, 15(2):8–16, August 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2815169.2815170.

Guido:1996:EIS

[GS96] Cardino Guido and Matina Skourti. Experiencing the introduction of software reused in Hitec. ACM SIGAPP Applied Computing Review, 4(2):15–20, September 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/251560.251583.

Gilesh:2018:SSV

[GSKJ18] M. P. Gilesh, Sanjay Satheesh, S. D. Madhu Kumar, and Lillykutty Jacob. Selecting suitable virtual machine migrations for optimal provisioning of virtual data centers. ACM SIGAPP Applied Computing Review, 18(2):22–32, July 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3243064.3243066.

Grambow:2020:BPM

[GWB20] Martin Grambow, Erik Wittern, and David Bermbach. Benchmarking the performance of microservice applications. ACM SIGAPP Applied Computing Review, 20(3):20–34, September 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3429204.
3429206.

Hoshino:2021:PAI

[HAG21] Shinji Hoshino, Yoshitaka Arahori, and Katsuhiko Gondow. Postmortem accurate IR-level state recovery for deployed concurrent programs. ACM SIGAPP Applied Computing Review, 21(3):33–48, September 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3493499.3493502.

Hansen:1995:HIS

[Han95] Steven C. Hansen. Hybrid inferential security methods for statistical databases. ACM SIGAPP Applied Computing Review, 3(1):14–18, June 1995. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/214310.214433.

Ho:2015:EHR

[HCC^{+}15] Chien-Chung Ho, Sheng-Wei Cheng, Yuan-Hao Chang, Yu-Ming Chang, Sheng-Yen Hong, and Che-Wei Chang. Efficient hibernation resuming with classification-based prefetching scheme for embedded computing systems. ACM SIGAPP Applied Computing Review, 15(1):33–43, March 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2753060.2753064.

Ho:2017:WAM

[HCC^{+}17] Chien-Chung Ho, Yu-Ming Chang, Yuan-Hao Chang, Hsiu-Chang Chen, and Tei-Wei Kuo. Write-aware memory management for hybrid SLC-MLC PCM memory systems. ACM SIGAPP Applied Computing Review, 17(2):16–26, August 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3131080.3131082.

Huang:2013:MBH

[HCT^{+}13] Po-Chun Huang, Yuan-Hao Chang, Che-Wei Tsao, Ming-Chang Yang, and Cheng-Kang Hsieh. Migration-based hybrid cache design for file systems over flash storage devices. ACM SIGAPP Applied Computing Review, 13(4):8–16, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2577554.2577556.

Huang:2014:GEE

[HDC^{+}14] Jun Huang, Qiang Duan, Qianbin Chen, Yu Sun, Yoshiaki Tanaka, and Wei Wang. Guaranteeing end-to-end quality-of-service with a generic routing approach. ACM SIGAPP Applied Computing Review, 14(2):8–22, June 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2656864.2656865.

Heckeler:2013:AMB

[HEK^{+}13] Patrick Heckeler, Hanno Eichelberger, Thomas Kropf, Jürgen Ruf, Stefan Huster, Sebastian Burg, Wolfgang Rosenstiel, and Bastian Schlich. Accelerated model-based robustness testing of state machine implementations. ACM SIGAPP Applied Computing Review, 13(3):50–67, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2537728.2537733.

Haas:2019:ACP

[HF19] Steffen Haas and Mathias Fischer. On the alert correlation process for the detection of multi-step attacks and a graphbased realization. ACM SIGAPP Applied Computing Review, 19(1):5–19, April 2019. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3325061.3325062.

Hwang:2015:PRB

[HK15] Won-Seok Hwang and Sang-Wook Kim. Post ranking in a blogosphere: algorithms and evaluation. ACM SIGAPP Applied Computing Review, 15(1):26–32, March 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2753060.2753063.

Han:2017:EMM

[HK17] Youngsub Han and Yanggon Kim. An extracting method of movie genre similarity using aspect-based approach in social media. ACM SIGAPP Applied Computing Review, 17(2):36– 45, August 2017. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3131080.3131084.

Hong:2019:ETM

[HLF^{+}19] Ding-Yong Hong, Shih-Kai Lin, Sheng-Yu Fu, Jan-Jan Wu, and Wei-Chung Hsu. Enhancing transactional memory execution via dynamic binary translation. ACM SIGAPP Applied Computing Review, 19(1):48–58, April 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3325061.3325065.

Hu:2012:UFD

[HMB12] Yeming Hu, Evangelos E. Milios, and James Blustein. A unified framework for document clustering with dual supervision. ACM SIGAPP Applied Computing Review, 12(2):53–63, June 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2340416.2340421.

Hagler:1995:CRW

[HMY95] Marion Hagler, William M. Marcy, and Jerry R. Yeargan. CD-ROM/WWW technology: a pragmatic approach to crossplatform archival publication. ACM SIGAPP Applied Computing Review, 3(2):16–19, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/228228.228234.

Huang:1993:ATS

[Hua93] Xiaoqiu Huang. Alignment of three sequences in quadratic space. ACM SIGAPP Applied Computing Review, 1(2):7–11, September 1993. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/381771.381773.

Hammerle-Uhl:2011:RWI

[HURU11] Jutta Hämmerle-Uhl, Karl Raab, and Andreas Uhl. Robust watermarking in iris recognition: application scenarios and impact on recognition performance. ACM SIGAPP Applied Computing Review, 11(3):6–18, August 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2034594.2034595.

Hadi:2024:DCN

[HW24] Abir Mohammad Hadi and Kwanghee Won. Deep convolutional neural network compression based on the intrinsic dimension of the training data. ACM SIGAPP Applied Computing Review, 24(1):14–23, March 2024. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3663652.3663654.

Huang:2022:DMH

[HXGB22] Jun Huang, Cong-Cong Xing, Shuyang Gu, and Erich Baker. Drop Maslow’s Hammer or not: machine learning for resource management in D2D communications. ACM SIGAPP Applied Computing Review, 22(1):5–14, March 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3530043.3530044.

Hematian:2017:TCB

[HYL^{+}17] Amirshahram Hematian, Wei Yu, Chao Lu, David Griffith, and Nada Golmie. Towards clustering-based device-to-device communications for supporting applications. ACM SIGAPP Applied Computing Review, 17(1):35–48, May 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/3090058.3090063.

Ishikawa:1996:MAP

[Ish96] Yutaka Ishikawa. MPC++ approach to parallel computing environment. ACM SIGAPP Applied Computing Review, 4(1): 15–18, April 1996. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/240732.240738.

Islam:2021:WCW

[IZ21] Md Rakibul Islam and Minhaz F. Zibran. What changes in where?: an empirical study of bug-fixing change patterns. ACM SIGAPP Applied Computing Review, 20(4):18–34, January 2021. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3447332.3447334.

Jannarone:1993:CIP

[Jan93] Robert J. Jannarone. Concurrent information processing, I: an applications overview. ACM SIGAPP Applied Computing Review, 1(2):1–6, September 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/381771.381772.

Jung:2017:LSD

[JCP17] Sungbo Jung, Dar-Jen Chang, and Juw Won Park. Large scale document inversion using a multi-threaded computing system. ACM SIGAPP Applied Computing Review, 17(2):27–35, August 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3131080.3131083.

Jamison:1999:SDA

[JL99] Wilfred C. Jamison and Doug Lea. Scripting distributed agents. ACM SIGAPP Applied Computing Review, 7(1):18–22, April 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
570150.570156.

Jeremic:2016:IRW

[JPM16] Nikolaus Jeremic, Helge Parzyjegla, and Gero Mühl. Improving random write performance in homogeneous and heterogeneous erasure-coded drive arrays. ACM SIGAPP Applied Computing Review, 15(4):31–53, February 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/2893706.2893709.

Krone:1998:CAE

[KCDS98] Oliver Krone, Fabrice Chantemargue, Thierry Dagaeff, and Michael Schumacher. Coordinating autonomous entities with STL. ACM SIGAPP Applied Computing Review, 6(2):18–32, September 1998. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/abs/10.
1145/297114.297118.

Kesselman:1996:HPP

[Kes96] Carl Kesselman. High performance parallel and distributed computation in compositional CC++. ACM SIGAPP Applied Computing Review, 4(1):24–26, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/240732.240741.

Kopacsi:1996:RSE

[KGKK96] Sándor Kopácsi, Daniela Gavalocová, George L. Kovács, and Ildikó Kmecs. Reuse in a simulation environment for flexible manufacturing systems. ACM SIGAPP Applied Computing Review, 4(2):2–7, September 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/251560.251581.

Kuo:2011:EFD

[KHC^{+}11] Tei-Wei Kuo, Po-Chun Huang, Yuan-Hao Chang, Chia-Ling Ko, and Chih-Wen Hsueh. An efficient fault detection algorithm for NAND flash memory. ACM SIGAPP Applied Computing Review, 11(2):8–16, March 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/1964144.1964146.

Kim:2010:RRT

[KJMJ10] Jungwook Kim, Seong Tae Jhang, Seung-Jin Moon, and Chu Shik Jhon. Register-relocation: a thermal-aware renaming method for reducing temperature of a register file. ACM SIGAPP Applied Computing Review, 11(1):41–51, June 2010. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
1869687.1869691.

Kopacsi:1997:COK

[KK97] Sádor Kopácsi and George L. Kovács. Co-operative knowledge processing in software reuse. ACM SIGAPP Applied Computing Review, 5(1):21–26, June 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/270895.270898.

Khlif:2017:MSM

[KKEK17] Ilhem Khlif, Mohamed Hadj Kacem, Cédric Eichler, and Ahmed Hadj Kacem. A multi-scale modeling approach for systems of systems architectures. ACM SIGAPP Applied Computing Review, 17(3):17–26, November 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3161534.3161536.

Khan:2017:USP

[KKH^{+}17] Arif Ali Khan, Jacky Keung, Shahid Hussain, Mahmood Niazi, and Muhammad Manzoor Ilahi Tamimy. Understanding software process improvement in global software development: a theoretical framework of human factors. ACM SIGAPP Applied Computing Review, 17(2):5–15, August 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3131080.3131081.

Kuo:2015:TAE

[KL15] Chin-Fu Kuo and Yung-Feng Lu. Task assignment with energy efficiency considerations for non-DVS heterogeneous multiprocessor systems. ACM SIGAPP Applied Computing Review, 14 (4):8–18, January 2015. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2724928.2724929.

Kuo:2017:SAP

[KL17] Chin-Fu Kuo and Yung-Feng Lu. Scheduling algorithm for parallel real-time tasks on multiprocessor systems. ACM SIGAPP Applied Computing Review, 16(4):14–24, January 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3040575.3040577.

Karre:2020:UUE

[KMR20] Sai Anirudh Karre, Neeraj Mathur, and Y. Raghu Reddy. Understanding usability evaluation setup for VR products in industry: a review study. ACM SIGAPP Applied Computing Review, 19(4):17–27, January 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3381307.3381309.

Kalra:2018:OBF

[KP18] Sumit Kalra and T. V. Prabhakar. Ontology-based framework for internal-external quality trade-offs and tenant management in multi-tenant applications. ACM SIGAPP Applied Computing Review, 17(4):46–58, January 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3183628.3183632.

Kolb:2013:FAA

[KR13] Jens Kolb and Manfred Reichert. A flexible approach for abstracting and personalizing large business process models. ACM SIGAPP Applied Computing Review, 13(1):6–18, March 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2460136.2460137.

Kostler:2021:STE

[KRHH21] Johannes Köstler, Hans P. Reiser, Gerhard Habiger, and Franz J. Hauck. SmartStream: towards efficient Byzantine resilient data streaming through speculation and sharding. ACM SIGAPP Applied Computing Review, 21(3):19–32, September 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3493499.
3493501.

Kao:2018:PBH

[KTH18] Chih-Chen Kao, Liang-Chi Tseng, and Wei-Chung Hsu. A pipeline-based heterogeneous framework for efficient synthetic light field rendering. ACM SIGAPP Applied Computing Review, 18(1):19–29, April 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3212069.3212071.

Kuhlenkamp:2020:AOF

[KWBE20] Jörn Kuhlenkamp, Sebastian Werner, Maria C. Borges, and Dominik Ernst. All but one: FaaS platform elasticity revisited. ACM SIGAPP Applied Computing Review, 20(3):5–19, September 2020. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3429204.3429205.

Liu:2001:LSF

[LBKV01] Damon Liu, Mark Burgin, Walter Karplus, and Daniel Valentino. Large-scale flow field visualization for aneurysm treatment. ACM SIGAPP Applied Computing Review, 9(1):3– 7, April 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
570142.570145.

Leitao:2013:ABS

[LC13] Luís Leitão and Pável Calado. An automatic blocking strategy for XML duplicate detection. ACM SIGAPP Applied Computing Review, 13(2):42–53, June 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2505420.2505424.

Liang:2021:CCD

[LCC^{+}21] Yu-Pei Liang, Shuo-Han Chen, Yuan-Hao Chang, Yun-Fei Liu, Hsin-Wen Wei, and Wei-Kuan Shih. A cache consolidation design of MLC STT–RAM for energy efficiency enhancement on cyber-physical systems. ACM SIGAPP Applied Computing Review, 21(1):37–49, March 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3477133.3477136.

Lin:2012:SOD

[LCL12] Szu-Yin Lin, Kuo-Ming Chao, and Chi-Chun Lo. Serviceoriented dynamic data driven application systems to urban traffic management in resource-bounded environment. ACM SIGAPP Applied Computing Review, 12(1):35–49, April 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2188379.2188383.

Lazreg:2018:FFA

[LCM18] Sami Lazreg, Philippe Collet, and Sébastien Mosser. Functional feasibility analysis of variability-intensive data floworiented applications over highly-configurable platforms. ACM SIGAPP Applied Computing Review, 18(3):32–48, October 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3284971.3284975.

Liu:2020:EDT

[LDY20] Xing Liu, Christophe De Vaulx, and Jingling Yuan. Experiment design for teaching digital logic and computer organization principle. ACM SIGAPP Applied Computing Review, 20 (1):24–35, April 2020. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3392350.3392353.

Lea:1998:SRS

[Lea98] Suzanne M. Lea. Simulating remotely sensed images of shoaling waves. ACM SIGAPP Applied Computing Review, 6(1): 19–25, April 1998. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/297090.297100.

Lindemann:2019:DAC

[LF19] Jens Lindemann and Mathias Fischer. On the detection of applications in co-resident virtual machines via a memory deduplication side-channel. ACM SIGAPP Applied Computing Review, 18(4):31–46, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3307624.3307628.

Lee:2023:DLR

[LFT23] Junhee Lee, Flavius Frasincar, and Maria Mihaela Trusca. DIWS-LCR-Rot-hop++: a domain-independent word selector for cross-domain aspect-based sentiment classification. ACM SIGAPP Applied Computing Review, 23(3):19–31, September 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3626307.
3626309.

Li:2012:ATG

[LHO12] Menglin Li, Seamus Hill, and Colm O’Riordan. Analysis of a triploid genetic algorithm over deceptive and epistatic landscapes. ACM SIGAPP Applied Computing Review, 12(3):51– 59, September 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2387358.2387362.

Leach:1993:ESG

[LK93] Stephen P. Leach and Abraham Kandel. Expert systems in government: a look at the redistricting problem. ACM SIGAPP Applied Computing Review, 1(1):2–9, January 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/152535.152536.

Lu:2013:RFT

[LK13] Yung-Feng Lu and Chin-Fu Kuo. Robust and flexible tunnel management for secure private cloud. ACM SIGAPP Applied Computing Review, 13(1):41–50, March 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2460136.2460140.

Lightbody:2017:EVF

[LKH17] Peter Lightbody, Tomás Krajník, and Marc Hanheide. An efficient visual fiducial localisation system. ACM SIGAPP Applied Computing Review, 17(3):28–37, November 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/3161534.3161537.

Lu:2012:NKM

[LKP12] Yung-Feng Lu, Chin-Fu Kuo, and Ai-Chun Pang. A novel key management scheme for wireless embedded systems. ACM SIGAPP Applied Computing Review, 12(1):50–59, April 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2188379.2188384.

Lo:2016:OSS

[LLC16] Shiwu Lo, Hung-Yi Lin, and Zhengyuan Chen. Optimizing swap space for improving process response after system resume. ACM SIGAPP Applied Computing Review, 15(4):54– 61, February 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2893706.2893710.

Lai:1993:ADE

[LLMW93] Albert Lai, Eric Lo, Wing Cheong Man, and Kam-Fai Wong. The application development environment of the DECmpp 12000 massively parallel computer — an introduction. ACM SIGAPP Applied Computing Review, 1(2):24–30, September 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
381771.381776.

Li:2012:MHC

[LLS12] Xinkai Li, Chao Lu, and Jon A. Sjogren. A method for Hensel code overflow detection. ACM SIGAPP Applied Computing Review, 12(1):6–11, April 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2188379.2188380.

Lazarova-Molnar:2017:MCO

[LMLAK17] Sanja Lazarova-Molnar, Halldór Tór Logason, Peter Grønbæk Andersen, and Mikkel Baun Kjærgaard. Mobile crowdsourcing of occupant feedback in smart buildings. ACM SIGAPP Applied Computing Review, 17(1):5–14, May 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/3090058.3090060.

Li:2016:MHA

[LNLC16] Jiajia Li, Grace Ngai, Hong Va Leong, and Stephen C. F. Chan. Multimodal human attention detection for reading from facial expression, eye gaze, and mouse dynamics. ACM SIGAPP Applied Computing Review, 16(3):37–49, November 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3015297.3015301.

Lee:2023:IBF

[LOCD23] Yeonjung Lee, Mert Ozer, Steven R. Corman, and Hasan Davulcu. Identifying behavioral factors leading to differential polarization effects of adversarial botnets. ACM SIGAPP Applied Computing Review, 23(2):44–56, June 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3610409.3610412.

Leithner:2021:CCH

[LS21] Manuel Leithner and Dimitris E. Simos. CHIEv: concurrent hybrid analysis for crawling and modeling of web applications. ACM SIGAPP Applied Computing Review, 21(1):5–23, March 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3477133.
3477134.

Lencastre:2023:PAG

[LSPC23] Maria Lencastre, Daniel Silva, João Henrique C. Pimentel, and Jaelson Brelaz Castro. PRIUS: Applying gamification to user stories prioritization. ACM SIGAPP Applied Computing Review, 23(4):27–44, December 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/10.1145/3642964.3642967.

Laine:2012:XEW

[LSV12] Markku Laine, Denis Shestakov, and Petri Vuorimaa. XFormsDB: an extensible web application framework built upon declarative W3C standards. ACM SIGAPP Applied Computing Review, 12(3):37–50, September 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2387358.2387361.

Li:1995:SIT

[LU95] Renqi Li and E. A. Unger. Security issues with TCP/IP. ACM SIGAPP Applied Computing Review, 3(1):6–13, June 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
214310.214313.

Lunardi:2018:EFJ

[LV18] Willian Tessaro Lunardi and Holger Voos. An extended flexible job shop scheduling problem with parallel operations. ACM SIGAPP Applied Computing Review, 18(2):46–56, July 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3243064.3243068.

Liao:2012:MCA

[LW12] Jhih-Jian Liao and Chin-Hsien Wu. A multi-controller architecture for high-performance solid-state drives. ACM SIGAPP Applied Computing Review, 12(4):58–66, December 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2432546.2432551.

Liu:2012:NUB

[LWC^{+}12] Jia Liu, Weiqing Wang, Zhenyu Chen, Xingzhong Du, and Qi Qi. A novel user-based collaborative filtering method by inferring tag ratings. ACM SIGAPP Applied Computing Review, 12(4):48–57, December 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2432546.2432550.

Lu:2014:PGS

[LWK14] Yung-Feng Lu, Jun Wu, and Chin-Fu Kuo. A path generation scheme for real-time green Internet of Things. ACM SIGAPP Applied Computing Review, 14(2):45–58, June 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2656864.2656868.

Liu:2023:CAG

[LWS^{+}23] Xing Liu, Ruixi Wang, Cai Shi, Chengming Zou, and Wenjie Zhu. Computing acceleration to genome-wide association study based on CPU/FPGA heterogeneous system. ACM SIGAPP Applied Computing Review, 23(4):16–26, December 2023. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3642964.3642966.

Likitvivatanavong:2013:ERC

[LY13] Chavalit Likitvivatanavong and Roland H. C. Yap. Eliminating redundancy in CSPs through merging and subsumption of domain values. ACM SIGAPP Applied Computing Review, 13 (2):20–29, June 2013. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2505420.2505422.

Marcelloni:2000:IOO

[MA00] Francesco Marcelloni and Mehmet Aksit. Improving objectoriented methods by using fuzzy logic. ACM SIGAPP Applied Computing Review, 8(2):14–23, December 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/373975.373982.

Mugeni:2022:GBB

[MA22] John Bosco Mugeni and Toshiyuki Amagasa. A graph-based blocking approach for entity matching using contrastively learned embeddings. ACM SIGAPP Applied Computing Review, 22(4):37–46, December 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/10.1145/3584014.3584017.

Marcal:2016:UMM

[MAM^{+}16] Edgar Marçal, Rossana Andrade, Rosemeiry Melo, Windson Viana, and Eduardo Junqueira. Using mobile message to improve student participation in blended courses: a Brazilian case study. ACM SIGAPP Applied Computing Review, 16(3):15– 25, November 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3015297.3015299.

Menor:2012:PPP

[MBP12] Mark Menor, Kyungim Baek, and Guylaine Poisson. Probabilistic prediction of protein phosphorylation sites using classification relevance units machines. ACM SIGAPP Applied Computing Review, 12(4):8–20, December 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2432546.2432547.

Maia:2013:FPC

[MBR13] Jose E. Bessa Maia, Angelo Brayner, and Fernando Rodrigues. A framework for processing complex queries in wireless sensor networks. ACM SIGAPP Applied Computing Review, 13(2): 30–41, June 2013. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2505420.2505423.

Marconi:2017:MDA

[MBR17] Francesco Marconi, Marcello M. Bersani, and Matteo Rossi. A model-driven approach for the formal verification of stormbased streaming applications. ACM SIGAPP Applied Computing Review, 17(3):6–15, November 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3161534.3161535.

May:2022:CPU

[MCMS22] Michael J. May, Yaakov Cohen, Hovav Menachem, and Yogev Swisa. CEMDA: Preserving the user’s mental model on Android files using complex events. ACM SIGAPP Applied Computing Review, 22(4):24–36, December 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3584014.3584016.

Medina:2015:IAW

[MCP15] Jonathas Leontino Medina, Maria Istela Cagnin, and Débora Maria Barroso Paiva. Investigating accessibility on web-based maps. ACM SIGAPP Applied Computing Review, 15(2):17– 26, August 2015. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2815169.2815171.

Martins:2019:EIV

[MCY^{+}19] Diogo S. Martins, Bruna C. R. Cunha, Cristiane A. Yaguinuma, Isabela Zaine, and Maria da Graça C. Pimentel. Effects of interactive video annotations on students’ browsing behavior and perceived workload. ACM SIGAPP Applied Computing Review, 19(2):44–57, August 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3357385.3357389.

Mengel:1995:KW

[Men95] Susan A. Mengel. K12 and the Web. ACM SIGAPP Applied Computing Review, 3(2):20–24, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/228228.228235.

Mao:2011:RLT

[MH11] Ziqing Mao and Cormac Herley. A robust link-translating proxy server mirroring the whole web. ACM SIGAPP Applied Computing Review, 11(2):30–42, March 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/1964144.1964149.

Meyer:1993:TFT

[MKR93] Chris Meyer, Abraham Kandel, and Dewey Rundus. The triad of fuzzy theory. ACM SIGAPP Applied Computing Review, 1 (2):12–15, September 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/381771.381774.

Matos:2012:SED

[MMC12] Luís Matos, José Moreira, and Alexandre Carvalho. A spatiotemporal extension for dealing with moving objects with extent in Oracle 11g. ACM SIGAPP Applied Computing Review, 12(2):7–17, June 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2340416.2340417.

Marchetti:2002:UXM

[MMSV02] Andrea Marchetti, Fabrizio Megale, Enrico Seta, and Fabio Vitali. Using XML as a means to access legislative documents: Italian and foreign experiences. ACM SIGAPP Applied Computing Review, 10(1):54–62, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/568235.568246.

Madhavan:2018:ESS

[MN18] Manu Madhavan and Gopakumar Gopalakrishnan Nair. An effective sequence structure representation for long non-coding RNA identification and cancer association using machine learning methods. ACM SIGAPP Applied Computing Review, 18(3): 49–58, October 2018. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3284971.3284976.

Martinez-Prieto:2012:QRD

[MPFC12] Miguel A. Martínez-Prieto, Javier D. Fernández, and Rodrigo Cánovas. Querying RDF dictionaries in compressed space. ACM SIGAPP Applied Computing Review, 12(2):64–77, June 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2340416.2340422.

Maranzato:2010:FDR

[MPNdL10] Rafael Maranzato, Adriano Pereira, Marden Neubert, and Alair Pereira do Lago. Fraud detection in reputation systems in e-markets using logistic regression and stepwise optimization. ACM SIGAPP Applied Computing Review, 11(1):14–26, June 2010. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
1869687.1869689.

Musilek:2000:SCE

[MPSR00] Petr Musílek, Witold Pedrycz, Giancarlo Succi, and Marek Reformat. Software cost estimation with fuzzy models. ACM SIGAPP Applied Computing Review, 8(2):24–29, December 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
373975.373984.

Munteanu:2001:EIE

[MR01] Cristian Munteanu and Agostinho Rosa. Evolutionary image enhancement with user behavior modeling. ACM SIGAPP Applied Computing Review, 9(1):8–14, April 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/570142.570146.

Mondal:2012:ESC

[MRS12] Manishankar Mondal, Chanchal K. Roy, and Kevin A. Schneider. An empirical study on clone stability. ACM SIGAPP Applied Computing Review, 12(3):20–36, September 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2387358.2387360.

Mukhopadhyay:1999:DCD

[Muk99] Snehasis Mukhopadhyay. Distributed control and distributed computing. ACM SIGAPP Applied Computing Review, 7(1): 23–24, April 1999. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/570150.570157.

Martins:2011:AAI

[MVOdGCP11] Diogo Santana Martins, Didier Augusto Vega-Oliveros, and Maria da Graça Campos Pimentel. Automatic authoring of interactive multimedia documents via media-oriented operators. ACM SIGAPP Applied Computing Review, 11(4):26– 37, December 2011. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2107756.2107759.

Moulema:2014:EMB

[MYX^{+}14] Paul Moulema, Wei Yu, Guobin Xu, David Griffith, Nada Golmie, Chao Lu, and David Su. On effectiveness of mesh-based protocols for smart grid communication networks.

ACM SIGAPP Applied Computing Review, 14(2):59–70, June 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2656864.2656869.

Morandin:1997:OFB

[MZ97] Elisabetta Morandin and Alessandro Zorer. An OO framework in the bulk data transfer domain. ACM SIGAPP Applied Computing Review, 5(2):16–20, September 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/297075.297083.

Majumder:2014:MSA

[MZAS14] Akm Jahangir Alam Majumder, Ishmat Zerin, Sheikh Iqbal Ahamed, and Roger O. Smith. A multi-sensor approach for fall risk prediction and prevention in elderly. ACM SIGAPP Applied Computing Review, 14(1):41–52, March 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2600617.2600621.

Norton:1996:POO

[NDS96] Charles D. Norton, Viktor K. Decyk, and Boleslaw K. Szymanski. On parallel object oriented programming in Fortran 90. ACM SIGAPP Applied Computing Review, 4(1):27–31, April 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
240732.240742.

Nguyen:2015:SRF

[NHM15] Vinh Nguyen, Chih-Cheng Hung, and Xiang Ma. Super resolution face image based on locally linear embedding and local correlation. ACM SIGAPP Applied Computing Review, 15(1): 17–25, March 2015. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2753060.2753062.

Niccolucci:2002:XFH

[Nic02] Franco Niccolucci. XML and the future of humanities computing. ACM SIGAPP Applied Computing Review, 10(1):43–47, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
568235.568244.

Naous:2021:PBA

[NL21] Dana Naous and Christine Legner. A preference-based approach to incorporate the “voice of the customer” in massmarket software product design. ACM SIGAPP Applied Computing Review, 20(4):35–49, January 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/10.1145/3447332.3447335.

Nardini:2011:CHS

[NOVS11] Elena Nardini, Andrea Omicini, Mirko Viroli, and Michael I. Schumacher. Coordinating e-health systems with TuCSoN semantic tuple centres. ACM SIGAPP Applied Computing Review, 11(2):43–53, March 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/1964144.1964150.

Notander:2013:SDS

[NRH13] Jesper Pedersen Notander, Per Runeson, and Martin Höst. SimPal: a design study on a framework for flexible safetycritical software development. ACM SIGAPP Applied Computing Review, 13(4):17–29, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2577554.2577558.

Neves:2021:DBB

[NVP21] Francisco Neves, Ricardo Vilaça, and José Pereira. Detailed black-box monitoring of distributed systems. ACM SIGAPP Applied Computing Review, 21(1):24–36, March 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/10.1145/3477133.3477135.

Nygate:1994:EEE

[Nyg94] Yossi Nygate. Ecxpert: exploiting event correlation in telecommunications. ACM SIGAPP Applied Computing Review, 2(2): 2–10, September 1994. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/381777.381778.

Nishimura:1998:CVV

[NYKI98] Toshikazu Nishimura, Hirofumi Yamaki, Takaaki Komura, and Toru Ishida. Community viewer: visualizing community formation on personal digital assistants. ACM SIGAPP Applied Computing Review, 6(1):13–18, April 1998. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/297090.297097.

Obaidat:1993:PEL

[OD93] M. S. Obaidat and D. L. Donahue. A priority Ethernet LAN protocol and its performance. ACM SIGAPP Applied Computing Review, 1(2):16–23, September 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/381771.381775.

Oyamada:2013:DSP

[OKK13] Masafumi Oyamada, Hideyuki Kawashima, and Hiroyuki Kitagawa. Data stream processing with concurrency control. ACM SIGAPP Applied Computing Review, 13(2):54–65, June 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2505420.2505425.

Ossendrijver:2023:ALM

[OSG23] Rick Ossendrijver, Stephan Schroevers, and Clemens Grelck. Automating library migrations with error prone and refaster. ACM SIGAPP Applied Computing Review, 23(1):5–19, March 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3594264.
3594265.

Okada:2001:CEE

[OTYM01] Masaya Okada, Hiroyuki Tarumi, Tetsuhiko Yoshimura, and Kazuyuki Moriya. Collaborative environmental education using distributed virtual environment accessible from real and virtual worlds. ACM SIGAPP Applied Computing Review, 9 (1):15–21, April 2001. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/570142.570147.

Oliveira:2014:HSC

[OWB^{+}14] Daniela Oliveira, Nicholas Wetzel, Max Bucci, Jesus Navarro, Dean Sullivan, and Yier Jin. Hardware–software collaboration for secure coexistence with kernel extensions. ACM SIGAPP Applied Computing Review, 14(3):22–35, September 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2670967.2670969.

Paranjape:2001:GE

[Par01] Raman B. Paranjape. Guest editorial. ACM SIGAPP Applied Computing Review, 9(3):1, September 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/570132.570133.

Pauwels:2019:DAH

[PC19] Stephen Pauwels and Toon Calders. Detecting anomalies in hybrid business process logs. ACM SIGAPP Applied Computing Review, 19(2):18–30, August 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3357385.3357387.

Phungtua-eng:2023:EDB

[PeSNY23] Thanapol Phungtua-eng, Shigeyuki Sako, Yushi Nishikawa, and Yoshitaka Yamamoto. Elastic data binning: Time-series sketching for time-domain astrophysics analysis. ACM SIGAPP Applied Computing Review, 23(2):5–22, June 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3610409.3610410.

Provensi:2013:UFP

[PEVR13] Lucas Provensi, Frank Eliassen, Roman Vitenberg, and Romain Rouvoy. Using fuzzy policies to improve context interpretation in adaptive systems. ACM SIGAPP Applied Computing Review, 13(3):26–37, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2537728.2537731.

Palanca:2015:TPS

[PHJJ15] Javier Palanca, Stella Heras, Javier Jorge, and Vicente Julian. Towards persuasive social recommendation: knowledge model. ACM SIGAPP Applied Computing Review, 15(2):41–49, August 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2815169.2815173.

Pijani:2020:IAP

[PIR20] Bizhan Alipour Pijani, Abdessamad Imine, and Michaël Rusinowitch. Inferring attributes with picture metadata embeddings. ACM SIGAPP Applied Computing Review, 20(2):36– 45, July 2020. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3412816.3412819.

Pape:2017:RDS

[PKBH17] Tobias Pape, Vasily Kirilichev, Carl Friedrich Bolz, and Robert Hirschfeld. Record data structures in Racket: usage analysis and optimization. ACM SIGAPP Applied Computing Review, 16(4):25–37, January 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3040575.3040578.

Poibrenski:2021:MMP

[PKVM21] Atanas Poibrenski, Matthias Klusch, Igor Vozniak, and Christian Müller. Multimodal multi-pedestrian path prediction for autonomous cars. ACM SIGAPP Applied Computing Review, 20(4):5–17, January 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3447332.3447333.

Park:2013:LHD

[PND^{+}13] Dongchul Park, Young Jin Nam, Biplob Debnath, David H. C. Du, Youngkyun Kim, and Youngchul Kim. An on-line hot data identification for flash-based storage using sampling mechanism. ACM SIGAPP Applied Computing Review, 13(1):51–64, March 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2460136.2460141.

Prellwitz:2018:AID

[PPM18] Matthias Prellwitz, Helge Parzyjegla, and Gero Mühl. Adaptive information distribution for dynamic sets using multicast push and pull. ACM SIGAPP Applied Computing Review, 18(3): 19–31, October 2018. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3284971.3284974.

Pack:2017:CAD

[PSS17] Chulwoo Pack, Seong-Ho Son, and Sung Shin. Computer aided diagnosis with boosted learning for anomaly detection in microwave tomography. ACM SIGAPP Applied Computing Review, 17(3):39–47, November 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3161534.3161538.

Perilo:2024:NBT

[PVT24] Michel Perilo, George Valença, and Aldenir Telles. Non-binary and trans-inclusive AI: a catalogue of best practices for developing automatic gender recognition solutions. ACM SIGAPP Applied Computing Review, 24(2):55–70, June 2024. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/10.1145/3687251.3687255.

Perez:2018:DEP

[PZ18] Alfredo J. Perez and Sherali Zeadally. Design and evaluation of a privacy architecture for crowdsensing applications. ACM SIGAPP Applied Computing Review, 18(1):7–18, April 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3212069.3212070.

Poltronieri:2018:UEF

[PZBdBC18] Ildevana Poltronieri, Avelino Francisco Zorzo, Maicon Bernardino, and Marcia de Borba Campos. Usability evaluation framework for domain-specific language: a focus group study. ACM SIGAPP Applied Computing Review, 18(3):5–18, October 2018. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3284971.3284973.

Qiao:2013:RNR

[QHW^{+}13] Ying Qiao, QiNan Han, Hongan Wang, Xiang Li, Kang Zhong, Keming Zhang, Jian Liu, and Anxiang Guo. RTRS: a novel real-time reasoning system based on active rules. ACM SIGAPP Applied Computing Review, 13(2):66–76, June 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2505420.2505426.

Qamhieh:2015:SBE

[QM15] Manar Qamhieh and Serge Midonnet. Simulation-based evaluations of DAG scheduling in hard real-time multiprocessor systems. ACM SIGAPP Applied Computing Review, 14(4): 27–39, January 2015. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2724928.2724931.

Raidl:1999:MCP

[Rai99] Günther R. Raidl. The multiple container packing problem: a genetic algorithm approach with weighted codings. ACM SIGAPP Applied Computing Review, 7(2):22–31, March 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
335527.335530.

Raj:1999:ACF

[Raj99a] Rajendra K. Raj. The active collections framework. ACM SIGAPP Applied Computing Review, 7(1):9–13, April 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/abs/10.1145/570150.570154.

Raje:1999:GEI

[Raj99b] Rajeev Raje. Guest editor’s introduction: distributed computing: a choice of the present and the future! ACM SIGAPP Applied Computing Review, 7(1):4, April 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/570150.570152.

Raje:1999:SDS

[Raj99c] Rajeev Raje. Session details: Special issue on distributed computing. ACM SIGAPP Applied Computing Review, 7(1):??, Spring 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3263630.

Raje:1999:DDC

[RC99] Rajeev R. Raje and Sivakumar Chinnasamy. Designing a distributed computing environment for global-scale systems: challenges and issues. ACM SIGAPP Applied Computing Review, 7(1):25–30, April 1999. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/570150.570158.

Ruy:2015:IBS

[RFB^{+}15] Fabiano B. Ruy, Ricardo A. Falbo, Monalessa P. Barcellos, Giancarlo Guizzardi, and Glaice K. S. Quirino. An ISO-based software process ontology pattern language and its application for harmonizing standards. ACM SIGAPP Applied Computing Review, 15(2):27–40, August 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2815169.2815172.

Rahman:2011:PPW

[RHA11] Farzana Rahman, Endadul Hoque, and Sheikh Iqbal Ahamed. Preserving privacy in wireless sensor networks using reliable data aggregation. ACM SIGAPP Applied Computing Review, 11(3):52–62, August 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2034594.2034599.

Rinaldi:2000:TCO

[Rin00] Andrea Rinaldi. Transforming a company with OOP. ACM SIGAPP Applied Computing Review, 8(1):20–25, September 2000. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
361651.361655.

Reis:2012:TIR

[RM12] Diogo Reis and Hugo Miranda. Transparently increasing RMI fault tolerance. ACM SIGAPP Applied Computing Review, 12 (2):18–26, June 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2340416.2340418.

Rosenblueth:1998:PPD

[Ros98] David A. Rosenblueth. A Prolog program for decomposing clothing orders into lays. ACM SIGAPP Applied Computing Review, 6(1):26–32, April 1998. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/297090.297112.

Rosenblueth:2019:GAD

[Ros19] David A. Rosenblueth. A greedy algorithm for decomposing large clothing orders into lays. ACM SIGAPP Applied Computing Review, 19(3):45–51, November 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/3372001.3372005.

Rauf:2011:DLB

[RP11] Irum Rauf and Ivan Porres. Designing level 3 behavioral RESTful web service interfaces. ACM SIGAPP Applied Computing Review, 11(3):19–31, August 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/2034594.2034596.

Rahman:2020:EKG

[RT20] Md Mostafizur Rahman and Atsuhiro Takasu. Exploiting knowledge graph and text for ranking entity types. ACM SIGAPP Applied Computing Review, 20(3):35–46, September 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3429204.
3429207.

Raje:2001:CSD

[RZW01] Rajeev R. Raje, Ming Zhong, and Tongyu Wang. Case study: a distributed concurrent system with AspectJ. ACM SIGAPP Applied Computing Review, 9(2):17–23, July 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/512000.512004.

Saeeda:2023:ICC

[SAG23] Hina Saeeda, Muhammad Ovais Ahmad, and Tomas Gustavsson. Identifying and categorizing challenges in large-scale agile software development projects: Insights from two Swedish companies. ACM SIGAPP Applied Computing Review, 23(2): 23–43, June 2023. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3610409.3610411.

Scordino:2019:RTE

[SAL19] Claudio Scordino, Luca Abeni, and Juri Lelli. Real-time and energy efficiency in Linux: theory and practice. ACM SIGAPP Applied Computing Review, 18(4):18–30, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3307624.3307627.

Succi:1996:ARI

[SB96] Giancarlo Succi and Francesco Baruchelli. Analysing the return of investment of reuse. ACM SIGAPP Applied Computing Review, 4(2):21–25, September 1996. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/251560.251584.

Sporer:2016:ASD

[SB16] Harald Sporer and Eugen Brenner. An automotive E/E system domain-specific modelling approach with various tool support. ACM SIGAPP Applied Computing Review, 16(1):5–14, April 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2924715.2924716.

Sharma:2012:DLS

[SD12] Anuj Sharma and Shubhamoy Dey. A document-level sentiment analysis approach using artificial neural network and sentiment lexicons. ACM SIGAPP Applied Computing Review, 12(4):67– 75, December 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2432546.2432552.

Sharma:2013:BSB

[SD13] Anuj Sharma and Shubhamoy Dey. A boosted SVM based ensemble classifier for sentiment analysis of online reviews. ACM SIGAPP Applied Computing Review, 13(4):43–52, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2577554.2577560.

Sobiech:2015:HAS

[SER15] Fabian Sobiech, Beate Eilermann, and Andreas Rausch. A heuristic approach to solve the elementary sprint optimization problem for non-cross-functional teams in Scrum. ACM SIGAPP Applied Computing Review, 14(4):19–26, January 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2724928.2724930.

Stach:2022:ATP

[SGPM22] Christoph Stach, Clémentine Gritti, Dennis Przytarski, and Bernhard Mitschang. Assessment and treatment of privacy issues in blockchain systems. ACM SIGAPP Applied Computing Review, 22(3):5–24, September 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/10.1145/3570733.3570734.

Sinkala:2023:GSM

[SH23] Zipani Tom Sinkala and Sebastian Herold. Generating seed mappings for machine learning-based code-to-architecture mappers using InMap. ACM SIGAPP Applied Computing Review, 23(4):5–15, December 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/10.1145/3642964.3642965.

Sayed:2022:CRD

[SHP22] Mohammed Sayed, Jae Yeon Hwang, and Juw Won Park. Circular RNA detection and effect of sequence homology. ACM SIGAPP Applied Computing Review, 22(1):24–30, March 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3530043.3530046.

Shahriar:2013:BOP

[SHV13] Hossain Shahriar, Hisham M. Haddad, and Ishan Vaidya. Buffer overflow patching for C and C++ programs: rule-based approach. ACM SIGAPP Applied Computing Review, 13(2): 8–19, June 2013. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2505420.2505421.

Saab:2016:PS

[SKEC16] Farah Saab, Ayman Kayssi, Imad Elhajj, and Ali Chehab. Playing with Sybil. ACM SIGAPP Applied Computing Review, 16(2):16–25, August 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/2993231.2993233.

Stamatopoulos:1993:TAS

[SKH93] Panagiotis Stamatopoulos, Isambo Karali, and Constantin Halatsis. A tour advisory system using a logic programming approach. ACM SIGAPP Applied Computing Review, 1(1):18– 25, January 1993. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/152535.152539.

Succi:1999:RBA

[SL99] Giancarlo Succi and Eric Liu. A relations-based approach for simplifying metrices extraction. ACM SIGAPP Applied Computing Review, 7(3):27–32, September 1999. CODEN ????

ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/333630.333635.

Sharma:2022:HHM

[SM22] Yanshul Sharma and Sanjay Moulik. HEAT: a heterogeneous multicore real-time scheduler with efficient energy and temperature management. ACM SIGAPP Applied Computing Review, 22(2):34–43, June 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3558053.3558056.

Shin:2017:SVB

[SMK17] Jungpil Shin, Ken Maruyama, and Cheol Min Kim. Signature verification based on inter-stroke and intra-stroke information. ACM SIGAPP Applied Computing Review, 17(1):26–34, May 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3090058.3090062.

SzymAnski:1996:OOP

[SN96] Boleslaw K. Szymanski and Charles D. Norton. Object oriented programming in parallel scientific computing: an overview of the special issue. ACM SIGAPP Applied Computing Review, 4(1):3–4, April 1996. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/240732.570160.

Sourek:2016:DSM

[SP16] Gustav Sourek and Petr Posík. Dynamic system modeling of evolutionary algorithms. ACM SIGAPP Applied Computing Review, 15(4):19–30, February 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/2893706.2893708.

Smith:2001:ABA

[SPB01] K. Smith, R. Paranjape, and L. Benedicenti. Agent behavior and agent models in unregulated markets. ACM SIGAPP Applied Computing Review, 9(3):2–12, September 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/570132.570134.

Spooren:2019:UDM

[SPD^{+}19] Jan Spooren, Davy Preuveneers, Lieven Desmet, Peter Janssen, and Wouter Joosen. On the use of DGAs in malware: an everlasting competition of detection and evasion. ACM SIGAPP Applied Computing Review, 19(2):31–43, August 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/3357385.3357388.

Seok:2011:EPC

[SPPP11] Hyunchul Seok, Youngwoo Park, Ki-Woong Park, and Kyu Ho Park. Efficient page caching algorithm with prediction and migration for a hybrid main memory. ACM SIGAPP Applied Computing Review, 11(4):38–48, December 2011. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2107756.2107760.

Schwartz:1993:UPM

[SS93] David G. Schwartz and Leon S. Sterling. Using a Prolog meta-programming approach for a blackboard application. ACM SIGAPP Applied Computing Review, 1(1):26–34, January 1993. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
152535.152542.

Stroulia:2002:DAR

[SS02] Eleni Stroulia and Tarja Systä. Dynamic analysis for reverse engineering and program understanding. ACM SIGAPP Applied Computing Review, 10(1):8–17, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/568235.568237.

Siboni:2018:ASM

[SSE18] Shachar Siboni, Asaf Shabtai, and Yuval Elovici. An attack scenario and mitigation mechanism for enterprise BYOD environments. ACM SIGAPP Applied Computing Review, 18 (2):5–21, July 2018. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3243064.3243065.

Staudt:1999:CBQ

[SSQJ99] Martin Staudt, René Soiron, Christoph Quix, and Matthias Jarke. Cost-based query optimization for metadata repositories. ACM SIGAPP Applied Computing Review, 7(2):10–21, March 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
335527.335529.

Sadeghi:2021:SBA

[SSR21] Mersedeh Sadeghi, Luca Sartor, and Matteo Rossi. A semanticbased access control approach for systems of systems. ACM SIGAPP Applied Computing Review, 21(4):5–19, December 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (elec-

tronic). URL https://dl.acm.org/doi/10.1145/3512753.
3512754.

Souza:2011:TBA

[SSS^{+}11] Jefferson R. Souza, Daniel O. Sales, Patrick Y. Shinzato, Fernando S. Osorio, and Denis F. Wolf. Template-based autonomous navigation and obstacle avoidance in urban environments. ACM SIGAPP Applied Computing Review, 11(4): 49–59, December 2011. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2107756.2107761.

Storey:2002:CDR

[SSW02] Margaret-Anne D. Storey, Susan Elliott Sim, and Kenny Wong. A collaborative demonstration of reverse engineering tools. ACM SIGAPP Applied Computing Review, 10(1):18–25, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
568235.568238.

Sioutas:2012:SSQ

[STP^{+}12] Spyros Sioutas, Dimitrios Tsoumakos, Alexandros Panaretos, Giannis Tzimas, Ioannis Karydis, and Dimitrios Tsolis. SART: speeding up query processing in sensor networks with an autonomous range tree structure. ACM SIGAPP Applied Computing Review, 12(3):60–74, September 2012. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2387358.2387363.

Succi:1999:RLP

[Suc99] Giancarlo Succi. The renewed life of parsing tools. ACM SIGAPP Applied Computing Review, 7(3):2–3, September 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
333630.333631.

Succi:1997:FVC

[SUV97] Giancarlo Succi, Carl Uhrik, and Tullio Vernazza. A formal view to classification and retrieval mechanism for reusable objects. ACM SIGAPP Applied Computing Review, 5(1):27–32, June 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
270895.270899.

Succi:1999:AJD

[SW99] Giancarlo Succi and Raymond W. Wong. The application of JavaCC to develop a C/C++ preprocessor. ACM SIGAPP Applied Computing Review, 7(3):11–18, September 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/333630.333633.

Shih:2016:MPT

[SW16] Chi-Sheng Shih and Guan-Fan Wu. Multiple protocol transport network gateway for IoT systems. ACM SIGAPP Applied Computing Review, 15(4):7–18, February 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2893706.2893707.

Shih:2013:FSV

[SWH^{+}13] Chi-Sheng Shih, Jie-Wen Wei, Shih-Hao Hung, Joen Chen, and Norman Chang. Fairness scheduler for virtual machines on heterogeneous multi-core platforms. ACM SIGAPP Applied Computing Review, 13(1):28–40, March 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2460136.2460139.

Shin:2020:SEP

[SWS20] Sangwon Shin, Kwanghee Won, and Sung Shin. Size efficient preprocessed symmetric RSA for wireless body area network. ACM SIGAPP Applied Computing Review, 20(1):15–23, April 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3392350.3392352.

Toledo:2013:HCA

[TAdOD13] Claudio F. M. Toledo, Marcio S. Arantes, Renato R. R. de Oliveira, and Alexandre C. B. Delbem. A hybrid cGA applied to the MLCLSP with overtime. ACM SIGAPP Applied Computing Review, 13(3):7–16, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2537728.2537729.

Tester:1994:ESR

[TBC94] Brian Tester, Steve Baker, and Barry Crabtree. An expert system for the real time control of the UK telephony network. ACM SIGAPP Applied Computing Review, 2(2):11– 17, September 1994. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/381777.381779.

Tseng:2017:CLF

[TCPC17] Hsueh-Wen Tseng, Wan-Chi Chang, I-Hsuan Peng, and Pei-Shan Chen. A cross-layer flow schedule with dynamical grouping for mitigating larger-scale TCP incast. ACM SIGAPP Applied Computing Review, 17(1):15–25, May 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/3090058.3090061.

Trivedi:2014:IBF

[TD14] Shrawan Kumar Trivedi and Shubhamoy Dey. Interaction between feature subset selection techniques and machine learning classifiers for detecting unsolicited emails. ACM SIGAPP Applied Computing Review, 14(1):53–61, March 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2600617.2600622.

Tseng:2014:EGS

[TFSO14] Hsueh-Wen Tseng, Yao-Chung Fan, Shiann-Tsong Sheu, and Shaiu-Yi Ou. An effective grouping scheme for avoiding hidden node problem in IEEE 802.15.4-based wireless sensor networks. ACM SIGAPP Applied Computing Review, 14(1):30–40, March 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2600617.2600620.

Tseng:2019:ARF

[TH19] Liang-Chi Tseng and Wei-Chung Hsu. An adaptive rendering framework for efficient synthetic light field generation. ACM SIGAPP Applied Computing Review, 19(3):33–44, November 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3372001.3372004.

Tang:2014:GBP

[THW14] Haokun Tang, Jun Huang, and Wei Wang. A game based passive worm defense model for P2P networks. ACM SIGAPP Applied Computing Review, 14(1):20–29, March 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2600617.2600619.

Tang:2012:SBD

[TJ12] Xiaolong Tang and Jaakko Järvi. Summary-based data-flow analysis that understands regular composite objects and iterators. ACM SIGAPP Applied Computing Review, 12(4):36– 47, December 2012. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2432546.2432549.

Teoman:2022:TAL

[TK22] Huseyin Alper Teoman and Pinar Karagoz. Trust-aware location recommendation for user groups. ACM SIGAPP Applied Computing Review, 22(3):39–55, September 2022. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3570733.3570736.

Thomas:1994:DSS

[TO94] Muffy Thomas and Bowen Ormsby. On the design of side-stick controllers in fly-by-wire aircraft. ACM SIGAPP Applied Computing Review, 2(1):15–20, March 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/381766.381769.

Trentini:2002:JBF

[Tre02] Andrea Trentini. A Java-based framework to support computer-assisted creation of structured XML documents. ACM SIGAPP Applied Computing Review, 10(1):48–53, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
568235.568245.

Tsetse:2013:EEI

[TWK^{+}13] Anthony K. Tsetse, Alexander L. Wijesinha, Ramesh Karne, Alae Loukili, and Patrick Appiah-Kubi. An experimental evaluation of IP4–IPV6 IVI translation. ACM SIGAPP Applied Computing Review, 13(1):19–27, March 2013. CODEN ????

ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/2460136.2460138.

Ulrich:2023:RTF

[UBA^{+}23] Jirí Ulrich, Jan Blaha, Ahmad Alsayed, Tomás Roucek, Farshad Arvin, and Tomás Krajník. Real time fiducial marker localisation system with full 6 DOF pose estimation. ACM SIGAPP Applied Computing Review, 23(1):20–35, March 2023. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3594264.3594266.

Urovi:2013:SPS

[UOB^{+}13] Visara Urovi, Alex C. Olivieri, Stefano Bromuri, Nicoletta Fornara, and Michael I. Schumacher. A semantic publish-subscribe coordination framework for IHE based crosscommunity health record exchange. ACM SIGAPP Applied Computing Review, 13(3):38–49, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2537728.2537732.

Valerio:1997:SIE

[Val97] Andrea Valerio. Special issue on the effects of frameworks and patterns on software reuse. ACM SIGAPP Applied Computing Review, 5(2):2–3, September 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/297075.570159.

Van:2002:SBX

[Van02] Huu Le Van. A system based on XML for supporting the management of educational web sites. ACM SIGAPP Applied Computing Review, 10(1):37–42, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/568235.568243.

vanDeursen:2002:SAR

[vD02] Arie van Deursen. Software architecture recovery and modelling: [WCRE 2001 discussion forum report]. ACM SIGAPP Applied Computing Review, 10(1):4–7, April 2002. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/568235.568236.

Vuong:2001:SAD

[VF01] Son T. Vuong and Peng Fu. A security architecture and design for mobile intelligent agent systems. ACM SIGAPP Applied Computing Review, 9(3):21–30, September 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/570132.570136.

Vale:1994:SES

[VFMM94] Zita A. Vale, M. Fernanda Fernandes, A. Machado e Moura, and Albino Marques. SPARSE — an expert system for alarm processing and operator assistance in substations control centers. ACM SIGAPP Applied Computing Review, 2(2):18–26, September 1994. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/381777.381780.

Vilela:2016:RAS

[VGH^{+}16] Jessyka Vilela, Enyo Goncalves, Ana Carla Holanda, Jaelson Castro, and Bruno Figueiredo. A retrospective analysis of SAC requirements: engineering track. ACM SIGAPP Applied Computing Review, 16(2):26–41, August 2016. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2993231.2993234.

Veloso:2020:CSU

[VGM^{+}20] Bruno Veloso, João Gama, Carlos Martins, Raphael Espanha, and Raul Azevedo. A case study on using heavy-hitters in interconnect bypass fraud. ACM SIGAPP Applied Computing Review, 20(3):47–57, September 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/10.1145/3429204.3429208.

Varniab:2020:DMI

[VHS20] Mahsa Shokri Varniab, Chih-Cheng Hung, and Vahid Khalilzad Sharghi. Data mining and image analysis using genetic programming. ACM SIGAPP Applied Computing Review, 19(4): 40–49, January 2020. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3381307.3381311.

Villacis:1999:NUJ

[Vil99] Juan Villacis. A note on the use of Java in scientific computing. ACM SIGAPP Applied Computing Review, 7(1):14–17, April 1999. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
570150.570155.

Valcarce:2019:DBT

[VPB19] Daniel Valcarce, Javier Parapar, and Álvaro Barreiro. Document-based and term-based linear methods for pseudorelevance feedback. ACM SIGAPP Applied Computing Review, 18(4):5–17, January 2019. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3307624.3307626.

Valerio:1997:DAF

[VSF97] Andrea Valerio, Giancarlo Succi, and Massimo Fenaroli. Domain analysis and framework-based software development. ACM SIGAPP Applied Computing Review, 5(2):4–15, September 1997. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
297075.297081.

Versick:2013:PCE

[VWT13] Daniel Versick, Ingolf Waßmann, and Djamshid Tavangarian. Power consumption estimation of CPU and peripheral components in virtual machines. ACM SIGAPP Applied Computing Review, 13(3):17–25, September 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2537728.2537730.

Wiegand:1994:CLP

[WA94] Nancy K. Wiegand and Teresa M. Adams. Comparative logical and physical modeling in two OODBMSs. ACM SIGAPP Applied Computing Review, 2(2):27–31, September 1994. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/381777.381781.

Warkentin:1995:CAW

[War95] Merrill E. Warkentin. Competitive advantage on the World Wide Web: a webmaster’s guide. ACM SIGAPP Applied Computing Review, 3(2):25–32, October 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/228228.228236.

Walker:2020:CCI

[WC20] Andrew Walker and Tomas Cerny. On cloud computing infrastructure for existing code-clone detection algorithms. ACM SIGAPP Applied Computing Review, 20(1):5–14, April 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
3392350.3392351.

Warkentin:1995:LCH

[WCR95] Merrill E. Warkentin, Christopher K. Carlson, and Stephen R. Ruth. A low-cost, high-yield intelligent decision support system for Navy contract termination processing. ACM SIGAPP Applied Computing Review, 3(1):27–30, June 1995. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/214310.214436.

Walker:2020:OST

[WCS20] Andrew Walker, Tomas Cerny, and Eungee Song. Open-source tools and benchmarks for code-clone detection: past, present, and future trends. ACM SIGAPP Applied Computing Review, 19(4):28–39, January 2020. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/3381307.3381310.

Won:2021:DLB

[WdCS21] Kwanghee Won, Hyung do Choi, and Sung Shin. Deep learningbased semantic classification of EMF-related scientific literature. ACM SIGAPP Applied Computing Review, 21(2):48–56, June 2021. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3477127.3477131.

Wu:2019:DMM

[WHC19] Chin-Hsien Wu, Cheng-Wei Huang, and Chen-Yu Chang. A data management method for databases using hybrid storage systems. ACM SIGAPP Applied Computing Review, 19(1): 34–47, April 2019. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3325061.3325064.

Won:2022:DII

[WJdCS22] Kwanghee Won, Youngsun Jang, Hyung do Choi, and Sung Shin. Design and implementation of information extraction system for scientific literature using fine-tuned deep learning models. ACM SIGAPP Applied Computing Review, 22(1):31– 38, March 2022. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3530043.3530047.

Wang:2021:TAP

[WLK^{+}21] Chuan-Chi Wang, Ying-Chiao Liao, Ming-Chang Kao, Wen-Yew Liang, and Shih-Hao Hung. Toward accurate platformaware performance modeling for deep neural networks. ACM SIGAPP Applied Computing Review, 21(1):50–61, March 2021. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3477133.3477137.

Wang:2001:TDP

[WW01] David K. Wang and James K. Wang. Towards the distributed processing of mobile software agents. ACM SIGAPP Applied Computing Review, 9(2):2–5, July 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/512000.512001.

Weerasinghe:2022:EDL

[WZL^{+}22] Shakthi Weerasinghe, Arkady Zaslavsky, Seng W. Loke, Alexey Medvedev, and Amin Abken. Estimating the dynamic lifetime of transient context in near real-time for cost-efficient adaptive caching. ACM SIGAPP Applied Computing Review, 22(2):44– 58, June 2022. CODEN ???? ISSN 1559-6915 (print), 1931-

0161 (electronic). URL https://dl.acm.org/doi/10.1145/
3558053.3558057.

Xie:2022:SNS

[XDS22] Liuyue Xie, Tinglin Duan, and Kenji Shimada. SAGA-net: shape-assisted graph attention neural network for real-time pointcloud completion. ACM SIGAPP Applied Computing Review, 22(2):21–33, June 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3558053.3558055.

Yohannes:2022:MNE

[YA22] Hailemariam Mehari Yohannes and Toshiyuki Amagasa. A method of named entity recognition for tigrinya. ACM SIGAPP Applied Computing Review, 22(3):56–68, September 2022. CO-DEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic).

URL https://dl.acm.org/doi/10.1145/3570733.3570737.

Yu:2015:TSM

[YAG^{+}15] Wei Yu, Dou An, David Griffith, Qingyu Yang, and Guobin Xu. Towards statistical modeling and machine learning based energy usage forecasting in smart grid. ACM SIGAPP Applied Computing Review, 15(1):6–16, March 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL

https://dl.acm.org/doi/abs/10.1145/2753060.2753061.

Yazidi:2016:QED

[YH16] Anis Yazidi and Hugo Hammer. Quantile estimation in dynamic and stationary environments using the theory of stochastic learning. ACM SIGAPP Applied Computing Review, 16(1): 15–24, April 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2924715.2924717.

Yajima:2001:SCS

[YHTN01] Etsuko Yajima, Takahiro Hara, Masahiko Tsukamoto, and Shojiro Nishio. Scheduling and caching strategies for correlated data in push-based information systems. ACM SIGAPP Applied Computing Review, 9(1):22–28, April 2001. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https:/

/dl.acm.org/doi/abs/10.1145/570142.570148.

Yokoyama:2014:ASP

[YKdSM14] Roberto Sadao Yokoyama, Bruno Yuji Lino Kimura, and Edson dos Santos Moreira. An architecture for secure positioning in a UAV swarm using RSSI-based distance estimation. ACM SIGAPP Applied Computing Review, 14(2):36–44, June 2014. CODEN ???? ISSN 1559-6915 (print), 1931-0161

(electronic). URL https://dl.acm.org/doi/abs/10.1145/
2656864.2656867.

Yang:2024:FTC

[YLJ24] Shasha Yang, Ziming Liu, and Lianghao Ji. Finite-time consensus of second-order multi-agent systems via event-triggered impulsive control. ACM SIGAPP Applied Computing Review, 24(1):5–13, March 2024. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/10.
1145/3663652.3663653.

Yuan:2022:FGR

[YLY^{+}22] Xu Yuan, Qihang Lei, Shuo Yu, Chengchuan Xu, and Zhikui Chen. Fine-grained relational learning for few-shot knowledge graph completion. ACM SIGAPP Applied Computing Review, 22(3):25–38, September 2022. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/10.1145/3570733.3570735.

Yuan:2019:AFD

[YPB19] Lanqin Yuan, Bernhard Pfahringer, and Jean Paul Barddal. Addressing feature drift in data streams using iterative subset selection. ACM SIGAPP Applied Computing Review, 19(1): 20–33, April 2019. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/3325061.3325063.

Yan:2017:AAA

[YSCX17] Hua Yan, Yulei Sui, Shiping Chen, and Jingling Xue. AutoFix: an automated approach to memory leak fixing on value-flow slices for C programs. ACM SIGAPP Applied Computing Review, 16(4):38–50, January 2017. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.

org/doi/abs/10.1145/3040575.3040579.

Zhang:2016:SMC

[ZCG^{+}16] Hanlin Zhang, Yevgeniy Cole, Linqiang Ge, Sixiao Wei, Wei Yu, Chao Lu, Genshe Chen, Dan Shen, Erik Blasch, and Khanh D. Pham. ScanMe mobile: a cloud-based Android malware analysis service. ACM SIGAPP Applied Computing Review, 16(1): 36–49, April 2016. CODEN ???? ISSN 1559-6915 (print),

1931-0161 (electronic). URL https://dl.acm.org/doi/abs/
10.1145/2924715.2924719.

Zhao:2015:JEB

[ZPH^{+}15] Yanxiao Zhao, Jems Pradhan, Jun Huang, Yu Luo, and Lina Pu. Joint energy-and-bandwidth spectrum sensing with GNU radio and USRP. ACM SIGAPP Applied Computing Review, 14(4):40–49, January 2015. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.acm.org/

doi/abs/10.1145/2724928.2724932.

Zibran:2013:GIF

[ZSRS13] Minhaz F. Zibran, Ripon K. Saha, Chanchal K. Roy, and Kevin A. Schneider. Genealogical insights into the facts and fictions of clone removal. ACM SIGAPP Applied Computing Review, 13(4):30–42, December 2013. CODEN ???? ISSN 1559-6915 (print), 1931-0161 (electronic). URL https://dl.

acm.org/doi/abs/10.1145/2577554.2577559.